### v1

# Generalized Entropy Calibration for Inference with Partially Observed Data: A Unified Framework

Abstract
Missing data is an universal problem in statistics. We develop a unified framework for estimating parameters defined by general estimating equations under a missing-at-random (MAR) mechanism, based on generalized entropy calibration weighting. We construct weights by minimizing a convex entropy subject to (i) balancing constraints on a data-adaptive calibration function, estimated using flexible machine-learning predictors with cross-fitting, and (ii) a debiasing constraint involving the fitted propensity score (PS) model. The resulting estimator is doubly robust, remaining consistent if either the outcome regression (OR) or the PS model is correctly specified, and attains the semiparametric efficiency bound when both models are correctly specified. Our formulation encompasses classical inverse probability weighting (IPW) and augmented IPW (AIPW) as special cases and accommodates a broad class of entropy functions. We illustrate the versatility of the approach in three important settings: semi-supervised learning with unlabeled outcomes, regression analysis with missing covariates, and causal effect estimation in observational studies. Extensive simulation studies and real-data applications demonstrate that the proposed estimators achieve greater efficiency and numerical stability than existing methods. In particular, the proposed estimator outperforms the classical AIPW estimator under the OR model misspecification.
Keywords: Calibration estimation, doubly robust, information projection, selection bias.
1 Introduction
Missing and incomplete data arise routinely in medical research, the social sciences, economics, and many other empirical disciplines. When the observed units are not representative of the target population, naïve complete-case analysis can produce biased and inefficient inference. A central goal of the missing-data literature is therefore to develop consistent and efficient estimators under plausible assumptions about the missingness mechanism (Little and Rubin, 2019; Kim and Shao, 2021).
Many modern data-analytic tasks can be viewed through the lens of partially observed data. In semi-supervised learning, the challenge is to exploit a large pool of unlabeled covariates for efficiency while guarding against covariate-shift bias (Chapelle et al., 2010; Zhang et al., 2019). In causal inference, systematic covariate imbalance between treatment groups must be corrected through weighting or matching (Hainmueller, 2012; Li et al., 2018). In regression with missing covariates, complete-case analysis is generally biased under missing at random (MAR; Ibrahim et al., 2005; Han and Wang, 2013). All three settings share a common structure: a parameter defined by an estimating equation that cannot be solved directly because part of the data vector is unobserved.
Two broad strategies exist. Outcome regression (OR) imputes the missing components via a conditional expectation model. Propensity-score (PS) weighting constructs inverse-probability weights (IPW). Doubly robust (DR) methods such as augmented IPW (AIPW; Robins et al., 1994) combine both and yield consistent estimators if either model is correctly specified. However, as demonstrated by Kang et al. (2007), standard DR estimators can be highly unstable with extreme weights. Targeted minimum loss-based estimation (TMLE; van der Laan and Rose, 2011) provides an alternative DR approach via a targeting step, but relies on a substitution estimator and does not directly produce calibration weights for the observed units. In parallel, entropy-based calibration methods choose weights by minimizing a convex divergence subject to balancing constraints (Zubizarreta, 2015; Chan et al., 2016; Zhao, 2019; Tan, 2020), achieving improved stability. These are closely related to covariate balancing methods in causal inference (Fan et al., 2022; Wang and Zubizarreta, 2020; Chattopadhyay et al., 2020; Hirshberg and Wager, 2021; Ben-Michael et al., 2021). However, existing calibration formulations do not directly yield a general doubly robust framework for arbitrary estimating equations with partially observed data.
In this paper, we develop a unified framework for analyzing partially observed data under MAR by combining generalized entropy calibration with doubly robust estimation and modern machine learning. The calibration weights are obtained by minimizing a generalized entropy subject to two sets of constraints: a balancing constraint based on a calibration function that approximates the optimal augmentation term, and a debiasing constraint incorporating a working PS model. To flexibly approximate the optimal calibration function, we embed cross-fitted machine-learning predictions into the calibration step, following the prediction-powered (Angelopoulos et al., 2023) and double machine learning (Chernozhukov et al., 2018) literatures.
The present work builds on the generalized entropy calibration framework of Kwon et al. (2025), which was developed for estimating finite population totals under design-based inference with known auxiliary variables. We extend their framework in three key directions. First, the target of inference is generalized from population totals to solutions of estimating equations $\mathbb{E}\{U(\theta;Z)\}=0$ under arbitrary MAR structures, which introduces a $\theta$-dependent calibration function requiring profile optimization (Algorithm 1) with no counterpart in the survey setting. Second, we establish double robustness and show that the proposed estimator achieves a strictly smaller asymptotic variance than AIPW under outcome model misspecification (Corollary 3), results that have no analog in the design-based framework where consistency follows from the known sampling mechanism. This variance-reduction property distinguishes our approach from both TMLE and the regularized calibrated estimation of Tan (2020), neither of which provides a similar mechanism under outcome-model misspecification. Third, we develop a cross-fitting procedure that integrates machine-learning predictions into the calibration constraints, and demonstrate the unifying scope of the approach across causal inference, semi-supervised learning, and missing covariate problems.
The main contributions are as follows. First, we formulate semi-supervised learning, regression with missing covariates, and causal inference within a single calibration-weighting framework for general estimating equations under MAR. Second, we developed a doubly robust generalized entropy calibration estimator that is doubly robust and locally efficient. Third, we developed a prediction-powered calibration. Cross-fitted machine-learning predictions enable flexible nonparametric calibration while maintaining valid large-sample theory. Simulations and real-data applications demonstrate substantial efficiency gains and improved stability, particularly under OR model misspecification.
The remainder of the paper is organized as follows. Section 2 introduces the setup and reviews IPW and AIPW. Section 3 presents the proposed estimators. Section 4 establishes large-sample properties. Section 5 discusses computation. Section 6 illustrates applications to causal inference, semi-supervised learning, and missing covariates. Section 7 reports simulations, and Section 8 presents real-data applications. Some concluding remarks are made in Section 9. All the technical proofs are relegated to the supplementary material (SM).
2 Preliminaries and Existing Methods
Let $Z=(X^{\top},Y)^{\top}$ be a $(p+1)$-dimensional random vector where $X\in\mathbb{R}^{p}$ denotes covariates and $Y$ is the outcome variable. Suppose we observe independent and identically distributed copies $Z_{1},\ldots,Z_{N}$ of $Z$. For each unit $i$, decompose the full data vector as
$Z_{i}=\bigl(O_{i}^{\top},M_{i}^{\top}\bigr)^{\top},$ (2.1)
where $O_{i}$ is the always-observed subvector and $M_{i}$ is the potentially missing subvector. This decomposition encompasses several familiar settings:
  1. 1.
    Missing outcomes: $O_{i}=X_{i}$, $M_{i}=Y_{i}$;
  2. 2.
    Missing covariates: $O_{i}=Y_{i}$, $M_{i}=X_{i}$;
  3. 3.
    Partially missing covariates: $O_{i}=(X_{1i},Y_{i})^{\top}$, $M_{i}=X_{2i}$, where $X_{i}=(X_{1i}^{\top},X_{2i}^{\top})^{\top}$.
Each setting is developed in detail in Section 6.
Let $\delta_{i}=1$ if $M_{i}$ is observed and $\delta_{i}=0$ otherwise, and assume a missing-at-random (MAR) mechanism (Rubin, 1976): $\delta_{i}\perp M_{i}\mid O_{i}$. Let $\theta\in\mathbb{R}^{q}$ denote the parameter of interest, defined as the unique solution to the population estimating equation $\mathbb{E}\{U(\theta;Z)\}=0$, where $U(\theta;Z)$ is a $q\times 1$ estimating function. Without missing data, $\theta$ can be estimated by solving
$\hat{U}_{N}(\theta)\equiv\sum_{i=1}^{N}U(\theta;z_{i})=0.$ (2.2)
When $M_{i}$ is missing for some units, $U(\theta;z_{i})$ is not fully computable and (2.2) cannot be solved directly.
Under MAR, a common strategy is to model the response mechanism via a propensity-score (PS) model,
$P(\delta_{i}=1\mid O_{i})=\pi(O_{i};\phi),$ (2.3)
where $\phi$ is a finite-dimensional parameter estimated by maximum likelihood. The inverse-probability-weighted (IPW) estimator of $\theta$ solves
$\sum_{i=1}^{N}\frac{\delta_{i}}{\pi(O_{i};\hat{\phi})}\,U(\theta;z_{i})=0.$ (2.4)
[…]
To improve efficiency, Robins et al. (1994) proposed the augmented IPW (AIPW) estimator, which solves
$\sum_{i=1}^{N}\left[\frac{\delta_{i}\,U(\theta;z_{i})}{\pi(O_{i};\hat{\phi})}-\frac{\delta_{i}-\pi(O_{i};\hat{\phi})}{\pi(O_{i};\hat{\phi})}\,b(O_{i})\right]=0,$ (2.5)
where $b(O_{i})$ is an arbitrary function of $O_{i}$.
The choice of $b(O_{i})$ is of central importance. A particularly important target is
$b^{*}(\theta;O_{i})=\mathbb{E}\{U(\theta;Z_{i})\mid O_{i}\},$ (2.6)
the conditional expectation of the estimating function given $O_{i}$. If an OR model is specified for $\mathbb{E}\{U(\theta;Z_{i})\mid O_{i}\}$ and correctly estimated, then using (2.6) in the AIPW estimator yields an estimator that is approximately unbiased even when the PS model is misspecified. Consequently, the AIPW estimator enjoys a doubly robust property if either the PS model or the OR model is correctly specified, and locally efficient if both are correct (Robins et al., 1994).
However, the AIPW estimator relies on an explicit augmentation term $b(O_{i})$ and can be sensitive to extreme inverse-probability weights, motivating the calibration-based approach developed in Section 3.
3 Proposed Estimators
We propose an alternative to AIPW that constructs data-adaptive weights $\omega_{i}=\omega(O_{i})$ through a generalized entropy calibration procedure. Rather than using an explicit augmentation term, the method incorporates auxiliary information implicitly via calibration constraints, preserving double robustness while improving numerical stability.
We define the estimator $\hat{\theta}_{\omega}$ as the solution to the weighted estimating equation
$\sum_{i=1}^{N}\delta_{i}\,\hat{\omega}_{i}(\theta)\,U(\theta;z_{i})=0,$ (3.1)
where $\hat{\omega}_{i}(\theta)=\hat{\omega}(\theta;O_{i})$ are calibration weights that may depend on $\theta$, since the relevant auxiliary information for estimating $\theta$ may itself depend on $\theta$.
To determine $\hat{\omega}_{i}(\theta)$, we employ the generalized entropy calibration method of Kwon et al. (2025). Let $G\colon\nu\to\mathbb{R}$ be a strictly convex, differentiable function with derivative $g=G^{\prime}$. The calibration weights for the observed units ($\delta_{i}=1$) are obtained by solving
$\min_{\omega_{1},\ldots,\omega_{N}}\sum_{i=1}^{N}\delta_{i}\,G(\omega_{i}),$ (3.2)
subject to the balancing constraint
$\sum_{i=1}^{N}\delta_{i}\,\omega_{i}\,b(O_{i})=\sum_{i=1}^{N}b(O_{i}),$ (3.3)
and the debiasing constraint
$\sum_{i=1}^{N}\delta_{i}\,\omega_{i}\,g(\hat{\pi}_{i}^{-1})=\sum_{i=1}^{N}g(\hat{\pi}_{i}^{-1}),$ (3.4)
[…]
Intuitively, if we use $b^{*}(O)=\mathbb{E}\left\{U(\theta;Z)\mid O\right\}$ in the balancing constraint and the OR model used to construct $b^{*}(O)$ is correct, balancing removes the leading bias term even when the PS model is wrong; conversely, if the PS model is correct, the debiasing constraint ensures consistency even when the OR model is wrong. These properties are formalized in Section 4.
Different choices of the entropy function $G$ produce different weighting schemes; several examples are summarized in Table 1.
Table 1: Examples of generalized entropies $G(\omega)$, the corresponding calibration covariates $\hat{g}_{i}=g(\hat{\pi}_{i}^{-1})$ and $\hat{g}_{i}^{-1}$.
Entropy $G(\omega)$ $\hat{g}_{i}$ $\hat{g}_{i}^{-1}$ Domain
Squared loss $\omega^{2}/2$ $\hat{\pi}_{i}^{-1}$ $\hat{\pi}_{i}$ $(-\infty,\infty)$
Empirical likelihood $\log\omega$ $\hat{\pi}_{i}$ $1/\hat{\pi}_{i}$ $(0,\infty)$
Exponential tilting $\omega\,\log\omega-\omega$ $\log(\hat{\pi}^{-1}_{i})$ $1/\log(\hat{\pi}^{-1}_{i})$ $(0,\infty)$
Hellinger distance $\sqrt{\omega}$ $\sqrt{\hat{\pi}_{i}/2}$ $\sqrt{2/\hat{\pi}_{i}}$ $(0,\infty)$
3.1 Dual formulation
The constrained optimization (3.2)–(3.4) is solved efficiently through its dual. Introduce Lagrange multipliers $\lambda_{1}\in\mathbb{R}^{q}$ and $\lambda_{2}\in\mathbb{R}$ for constraints (3.3)–(3.4), and write $b_{i}=b(O_{i})$, $\hat{g}_{i}=g(\hat{\pi}_{i}^{-1})$, and $s_{i}=(b_{i}^{\top},\hat{g}_{i})^{\top}$. The Lagrangian is
$Q(\omega,\lambda)$ $=-\sum_{i=1}^{N}\delta_{i}\,G(\omega_{i})+\lambda_{1}^{\top}\left(\sum_{i=1}^{N}\delta_{i}\,\omega_{i}\,b_{i}-\sum_{i=1}^{N}b_{i}\right)+\lambda_{2}\left(\sum_{i=1}^{N}\delta_{i}\,\omega_{i}\,\hat{g}_{i}-\sum_{i=1}^{N}\hat{g}_{i}\right),$ (3.5)
where $\lambda=\left(\lambda_{1}^{\top},\lambda_{2}\right)^{\top}$. Maximizing $Q(\omega,\lambda)$ with respect to $\omega_{i}$ yields the closed-form optimal weights
$\omega_{i}^{*}(\lambda)=g^{-1}(\lambda_{1}^{\top}b_{i}+\lambda_{2}\hat{g}_{i})=g^{-1}(\lambda^{\top}s_{i}),$ (3.6)
where $\lambda=(\lambda_{1}^{\top},\lambda_{2})^{\top}$ and $g^{-1}$ is strictly increasing because $G$ is strictly convex. The function $g(\cdot)$ is called the calibration link function and is closely related to the canonical link function in the generalized linear models. The calibration link operates on the weight parameter $\omega_{i}$ rather than a conditional mean $\mu_{i}$, but the algebraic structure is identical. Substituting (3.6) back to (3.5) gives the dual objective
$\rho_{G}(\lambda)=\frac{1}{N}\sum_{i=1}^{N}\delta_{i}\,F(\lambda^{\top}s_{i})-\frac{1}{N}\sum_{i=1}^{N}\lambda^{\top}s_{i},$ (3.7)
where $F(\nu)=-G\{g^{-1}(\nu)\}+g^{-1}(\nu)\,\nu$ is the convex conjugate of $G$. The optimal dual solution
$\hat{\lambda}=\arg\min_{\lambda}\rho_{G}(\lambda)$
[…]
Differentiating (3.7) confirms that $\hat{\lambda}$ satisfies the calibration equations (3.3)–(3.4):
$\frac{\partial}{\partial\lambda}\rho_{G}(\lambda)=\frac{1}{N}\left\{\sum_{i=1}^{N}\delta_{i}\,\omega_{i}^{*}(\lambda)\,s_{i}-\sum_{i=1}^{N}s_{i}\right\}.$
For this reason, we refer to $\rho_{G}(\lambda)$ as the calibration generating function induced by the entropy $G$. The dual representation reduces the problem to optimizing over $\lambda\in\mathbb{R}^{q+1}$, regardless of the sample size $N$.
4 Large-Sample Properties
We establish the asymptotic properties of the proposed estimator $\hat{\theta}_{\omega}$ defined in (3.1). Throughout, the estimator solves
$\hat{U}^{\omega}(\theta)\;\equiv\;\sum_{i=1}^{N}\delta_{i}\,\omega_{i}^{*}(\hat{\lambda},\hat{\varphi})\,U(\theta;\,z_{i})\;=\;0,$
where $\omega_{i}^{*}(\lambda,\varphi)=g^{-1}\!\bigl(\lambda_{1}^{\top}b_{i}+\lambda_{2}\,g_{i}(\varphi)\bigr)$, $b_{i}=b(O_{i})$, and $g_{i}(\varphi)=g\{\pi^{-1}(O_{i};\varphi)\}$. The parameters $\hat{\varphi}$ and $\hat{\lambda}$ are obtained jointly from
$\nabla\ell(\varphi)$ $\;\equiv\;\frac{1}{N}\sum_{i=1}^{N}\biggl(\frac{\delta_{i}}{\pi_{i}}-1\biggr)h_{i}(\varphi)\;=\;0,$ (4.1)
$\nabla\hat{\rho}_{G}(\lambda)$ $\;\equiv\;\frac{1}{N}\sum_{i=1}^{N}\Biggl\{\delta_{i}\,\omega_{i}^{*}(\lambda,\hat{\varphi})\begin{pmatrix}b_{i}\\ g_{i}(\hat{\varphi})\end{pmatrix}-\begin{pmatrix}b_{i}\\ g_{i}(\hat{\varphi})\end{pmatrix}\Biggr\}\;=\;0,$ (4.2)
where $\pi_{i}=\pi(O_{i};\varphi)$ and $h_{i}(\varphi)=\{1-\pi_{i}(\varphi)\}^{-1}\,\partial\pi_{i}(\varphi)/\partial\varphi$.
The regularity conditions are collected in the Supplementary Material (SM), Section A. Assumptions 1–2 ensure the consistency of $\hat{\varphi}$ and $\hat{\lambda}$, established in Lemma S.1 of the SM. Assumptions 3–4 impose smoothness and non-degeneracy of the limiting dual and propensity-score objectives; in particular, non-singularity of hessian $\mathbb{E}\{\nabla^{2}\rho_{G}(\lambda^{*})\}$, where $\lambda^{*}=\arg\min_{\lambda}\rho_{G}(\lambda)$ and $\rho_{G}(\lambda)$ is defined in (3.7), requires that the calibration covariates $s_{i}=(b_{i}^{\top},g_{i})^{\top}$ are not collinear among the respondents, which is the natural identifiability condition for entropy calibration. Assumptions 5–7 are standard conditions for joint M-estimation of the system (4.1)–(4.2); they hold whenever the three sub-problems are individually well-posed and their interaction is smooth.
4.1 Linearization
The following theorem provides a first-order linear expansion of $\hat{U}^{\omega}(\theta)$ that underlies all subsequent asymptotic results.
Theorem 4.1 (Linearization).
Under Assumptions 1–4,
$\hat{U}^{\omega}(\theta)\;=\;\tilde{U}^{\omega}(\lambda^{*},\hat{\varphi})\;+\;o_{p}(N^{-1/2}),$
where
[…]
$\sum_{i=1}^{N}\delta_{i}\,f^{\prime}\!\bigl(\lambda^{*\top}s_{i}(\hat{\varphi})\bigr)\,\bigl\{U(\theta;\,z_{i})-\gamma\,s_{i}(\hat{\varphi})\bigr\}\,s_{i}^{\top}(\hat{\varphi})\;=\;0,$
with $f^{\prime}=(g^{-1})^{\prime}$. This expansion holds without assuming correctness of either the PS or the OR model.
[…]
4.2 Consistency under the correct PS model
We first show that the calibration weights reduce to IPW weights when the PS model is correct, and derive the resulting asymptotic distribution.
Lemma 4.1.
If the PS model is correctly specified, i.e., $P(\delta=1\mid O)=\pi(O;\varphi_{0})$, then $\varphi^{*}=\varphi_{0}$ and $\lambda_{1}^{*}\to 0$, $\lambda_{2}^{*}\to 1$, so that $\omega_{i}^{*}(\lambda^{*},\varphi_{0})=1/\pi(O_{i};\varphi_{0})$.
Corollary 1 (Asymptotic normality under correct PS model).
Under Assumptions 1–7 and a correctly specified PS model,
$\sqrt{N}\,(\hat{\theta}_{\omega}-\theta_{0})\;\xrightarrow{d}\;\mathcal{N}\!\bigl(0,\;\tau_{1}^{-1}\,V_{1}\,(\tau_{1}^{-1})^{\top}\bigr),$
where $\tau_{1}=\mathbb{E}\bigl\{\partial U(\theta_{0};Z)/\partial\theta^{\top}\bigr\}$ and
$V_{1}\;=\;\mathrm{Var}\bigl\{U(\theta_{0};Z)\bigr\}\;+\;\mathbb{E}\!\left[\left(\frac{1}{\pi(O;\varphi_{0})}-1\right)\left\{U(\theta_{0};Z)-\gamma^{*}S(O;\varphi_{0})-\kappa^{*}h(O;\varphi_{0})\right\}^{\otimes 2}\right],$
with $S(O;\varphi_{0})=\bigl(b^{\top}(O),\,g(\pi^{-1}(O;\varphi_{0}))\bigr)^{\top}$ and $B^{\otimes}=BB^{\top}$. Here $\kappa^{*}\in\mathbb{R}^{q\times r}$ is the probability limit of $\hat{\kappa}$ defined by
$\frac{1}{N}\sum_{i=1}^{N}\frac{\partial}{\partial\varphi}\biggl\{\gamma^{*}s_{i}(\varphi)+\delta_{i}\,\omega_{i}^{*}(\lambda^{*},\varphi)\bigl[U(\theta;\,z_{i})-\gamma^{*}s_{i}(\varphi)\bigr]+\Bigl(1-\frac{\delta_{i}}{\pi(O_{i};\varphi)}\Bigr)\kappa\,h_{i}(\varphi)\biggr\}=0.$ (4.3)
The term $\kappa^{*}h(O;\varphi_{0})$ captures the effect of estimating $\varphi$ on the asymptotic variance; it vanishes when the OR model is also correctly specified.
4.3 Consistency under the correct OR model
Lemma 4.2.
If the OR model is correctly specified, i.e., $\mathbb{E}\{U(\theta_{0};Z)\mid O\}\in\mathrm{span}\{b(O)\}$, and $b^{*}(O)=\mathbb{E}\{U(\theta;Z)\mid O\}$ is used in (3.3), then $\kappa^{*}=0$.
Corollary 2 (Asymptotic normality under correct OR model).
Under Assumptions 1–7, if $b^{*}(O)=\mathbb{E}\{U(\theta;Z)\mid O\}$ satisfies $\mathbb{E}\{U(\theta_{0};Z)\mid O\}\in\mathrm{span}\{b(O)\}$, then
$\sqrt{N}\,(\hat{\theta}_{\omega}-\theta_{0})\;\xrightarrow{d}\;\mathcal{N}\!\bigl(0,\;\tau_{1}^{-1}\,\bar{V}_{1}\,(\tau_{1}^{-1})^{\top}\bigr),$
where
$\bar{V}_{1}\;=\;\mathrm{Var}\bigl\{\mathbb{E}[U(\theta_{0};Z)\mid O]\bigr\}\;+\;\mathbb{E}\!\Big[\delta\,\bigl\{\omega^{*}(O;\lambda^{*},\varphi^{*})\bigr\}^{2}\,\mathrm{Var}\{U(\theta_{0};Z)\mid O\}\Big].$
Hence $\hat{\theta}_{\omega}$ remains $\sqrt{N}$-consistent and asymptotically normal even when the PS model is misspecified.
4.4 Double robustness, variance dominance, and local efficiency
Corollaries 1 and 2 together establish that $\hat{\theta}_{\omega}$ is doubly robust: it is consistent for $\theta_{0}$ if either the PS or the OR model is correctly specified. We now show that, under a correct PS model, the proposed estimator is never less efficient than the classical AIPW estimator and is strictly more efficient whenever the OR model is misspecified and the debiasing covariate carries additional information.
Corollary 3 (Variance dominance over AIPW).
Under the conditions of Corollary 1, let $V_{1}$ denote the variance component of the proposed estimator $\hat{\theta}_{\omega}$ and let
$V_{3}\;=\;\mathrm{Var}\bigl\{U(\theta_{0};Z)\bigr\}\;+\;\mathbb{E}\!\left[\left(\frac{1}{\pi(O;\varphi_{0})}-1\right)\left\{U(\theta_{0};Z)-b(O)\right\}^{\otimes 2}\right]$
be the corresponding variance component of the classical AIPW estimator. Then
$V_{1}\;\leq\;V_{3},$ (4.4)
with equality if and only if $g\bigl(\pi^{-1}(O;\varphi_{0})\bigr)$ and $h(O;\varphi_{0})$ lie in $\mathrm{span}\{b(O)\}$ almost surely.
The variance reduction is most pronounced when the OR model is substantially misspecified, so that $b(O)$ is a poor approximation to $\mathbb{E}\{U(\theta_{0};Z)\mid O\}$, but the PS model provides useful information about the missingness mechanism through the debiasing covariate $g(\pi^{-1}(O;\varphi_{0}))$ and the score function $h(O;\varphi_{0})$. In such settings, the entropy calibration framework effectively enriches the augmentation space beyond what AIPW uses, yielding strictly smaller asymptotic variance without requiring any additional modeling effort.
Remark 1.
When both models are correctly specified and $b^{*}(O)=\mathbb{E}\{U(\theta;Z)\mid O\}$ is used in calibration, both $V_{1}$ and $V_{3}$ reduce to
$V_{1}^{*}\;=\;\mathrm{Var}\bigl\{U(\theta_{0};Z)\bigr\}\;+\;\mathbb{E}\!\left[\left(\frac{1}{\pi(O;\varphi_{0})}-1\right)\mathrm{Var}\{U(\theta_{0};Z)\mid O\}\right],$
which coincides with the semiparametric efficiency lower bound of Robins et al. (1994). Thus, the proposed estimator is also locally efficient.
5 Computational Details
This section describes the numerical implementation of $\hat{\theta}_{\omega}$. The dual formulation in Section 3.1 reduces the calibration step to an unconstrained convex minimization over $\lambda\in\mathbb{R}^{q+1}$. We now address two additional computational issues: (i) solving the nested optimization when the calibration function $b(O_{i})$ depends on $\theta$, and (ii) constructing $b(O_{i})$ via cross-fitted machine-learning predictions.
5.1 Profile optimization
The optimal calibration function is $b^{*}(\theta;O_{i})=\mathbb{E}\{U(\theta;Z_{i})\mid O_{i}\}$, which depends on the unknown parameter $\theta$. Under this choice, the dual objective becomes $\rho_{G}(\lambda,\theta)$, with $)s_{i}(\theta)=(b^{*\top}(\theta;O_{i}),\,g_{i}(\hat{\phi}$$)^{\top}$, and the estimator solves the saddle-point problem
$\hat{\theta}_{\omega}=\arg\max_{\theta}\min_{\lambda}\;\rho_{G}(\lambda,\theta).$ (5.1)
Equivalently, defining the profile objective $L(\theta)=\rho_{G}(\hat{\lambda}(\theta),\theta)$ where $\hat{\lambda}(\theta)=\arg\min_{\lambda}\rho_{G}(\lambda,\theta)$, we have $\hat{\theta}_{\omega}=\arg\max_{\theta}L(\theta)$.
We solve (5.1) using a two-loop procedure. For a fixed $\theta$, the inner loop minimizes $\rho_{G}(\lambda,\theta)$ over $\lambda$ using Newton’s method. The outer loop updates $\theta$ via a Newton or BFGS step based on the gradient of $L(\theta)$. By the envelope theorem (Danskin, 2012), $\nabla_{\theta}L(\theta)=\nabla_{\theta}\rho_{G}(\hat{\lambda}(\theta),\theta)$, since $\nabla_{\lambda}\rho_{G}=0$ at the inner optimum. The procedure is summarized in Algorithm 1.
Algorithm 1 Two-Loop Profile Optimization for $\hat{\theta}_{\omega}$
1:Initial value $\theta^{(0)}$; tolerances $\epsilon>0$, $\delta>0$
[…]
3:repeat
4:  Inner loop: Compute $\hat{\lambda}(\theta^{(k)})=\arg\min_{\lambda}\rho_{G}(\lambda,\theta^{(k)})$
5:  Outer loop: Update $\theta^{(k+1)}\leftarrow\theta^{(k)}-[\nabla^{2}_{\theta}L(\theta^{(k)})]^{-1}\,\nabla_{\theta}L(\theta^{(k)})$
6:  $k\leftarrow k+1$
7:until $\|\theta^{(k)}-\theta^{(k-1)}\|\leq\epsilon$ or $\|\nabla_{\theta}L(\theta^{(k-1)})\|\leq\delta$
8:return $\hat{\theta}_{\omega}\leftarrow\theta^{(k)}$
5.2 Calibration via cross-fitting
In practice, the conditional expectation $b^{*}(\theta;O_{i})=\mathbb{E}\{U(\theta;Z_{i})\mid O_{i}\}$ is unknown and must be estimated. Following the idea of Angelopoulos et al. (2023), we approximate it by plugging in a prediction $\hat{M}_{i}$ of the missing component $M_{i}$: $b^{*}(\theta;O_{i})\approx U(\theta;\,O_{i},\hat{M}_{i})$, where $\hat{M}_{i}$ is obtained from a flexible machine-learning model fitted to the observed data. To avoid the overfitting bias that arises from using the same data to train the prediction rule and to estimate $\theta$ (Chernozhukov et al., 2018), we adopt $K$-fold cross-fitting. Let $\mathcal{I}=\{1,\ldots,N\}$ and $\mathcal{S}=\{i\in\mathcal{I}\mathrel{\mathop{\ordinarycolon}}\delta_{i}=1\}$. We randomly partition $\mathcal{I}$ into $K$ disjoint folds $\{\mathcal{I}^{(k)}\}_{k=1}^{K}$. For each fold $k$, we fit a prediction model $\hat{m}^{(-k)}(\cdot)$ using training data $\mathcal{S}^{(-k)}=\mathcal{S}\setminus(\mathcal{S}\cap\mathcal{I}^{(k)})$ and compute out-of-fold predictions $\hat{M}_{i}^{(-k)}=\hat{m}^{(-k)}(O_{i})$ for $i\in\mathcal{I}^{(k)}$. Using these cross-fitted predictions, we solve the calibration problem:
$\min_{\omega_{1},\ldots,\omega_{N}}\sum_{i=1}^{N}\delta_{i}\,G(\omega_{i}),$ (5.2)
$\text{subject to}\quad\sum_{k=1}^{K}\sum_{i\in\mathcal{I}^{(k)}}\delta_{i}\,\omega_{i}\,U\!\left(\theta;\,O_{i},\hat{M}_{i}^{(-k)}\right)=\sum_{k=1}^{K}\sum_{i\in\mathcal{I}^{(k)}}U\!\left(\theta;\,O_{i},\hat{M}_{i}^{(-k)}\right),$ (5.3)
[…]
The two-loop procedure in Algorithm 1 is then applied to obtain $\hat{\theta}_{\omega}$.
Note that the prediction model $\hat{m}^{(-k)}$ is trained only on the complete cases in $\mathcal{S}^{(-k)}$, but the out-of-fold predictions $\hat{M}_{i}^{(-k)}$ are computed for all units $i\in\mathcal{I}^{(k)}$, including those with $\delta_{i}=0$. This is essential: the balancing constraint (5.3) involves the full sample, so calibration function values are needed for every unit regardless of whether $M_{i}$ is observed.
Remark 2.
The asymptotic results in Section 4 require the cross-fitted predictor $\hat{m}^{(-k)}$ to satisfy a mean-squared-error condition of the form $\mathbb{E}\|\hat{M}_{i}^{(-k)}-\mathbb{E}(M_{i}\mid O_{i})\|^{2}=o(1)$, i.e., the prediction error must vanish in probability. This condition is mild and is satisfied by a broad range of modern machine-learning methods, including random forests, gradient boosting, neural networks, and penalized regression, provided the sample size grows and standard regularity conditions hold. No specific convergence rate (e.g., $N^{-1/4}$) is required for the cross-fitted predictions, because the debiasing constraint ensures that the estimator remains consistent even when the prediction model converges slowly, as long as the PS model is correctly specified. When neither model is correctly specified, faster convergence of the predictor generally improves finite-sample performance, though consistency is no longer guaranteed. See Section C of SM for more details.
6 Illustrative Examples
We now describe how the proposed framework applies to three important settings. In each case, we specify the decomposition $(O_{i},M_{i})$, the estimating function $U(\theta;Z_{i})$, and the calibration function $b(\theta;O_{i})$; the calibration weights and estimator are then obtained by applying the general procedure in Sections 35.
6.1 Causal inference
Let $T_{i}\in\{0,1\}$ be a binary treatment indicator, $Y_{i}(1)$ and $Y_{i}(0)$ are the potential outcomes, and $X_{i}\in\mathbb{R}^{p}$ the observed covariates. Under SUTVA, unconfoundedness, and positivity (Rubin, 1974; Rosenbaum and Rubin, 1983; Imbens and Rubin, 2015), the observed outcome is $Y_{i}=T_{i}Y_{i}(1)+(1-T_{i})Y_{i}(0)$ and the average treatment effect (ATE) is
$\theta=\theta_{1}-\theta_{0}=\mathbb{E}\{Y(1)\}-\mathbb{E}\{Y(0)\},$
where $\theta_{t}$ solves $\mathbb{E}\{U_{t}(\theta_{t};X,Y(t))\}=0$ with $U_{t}(\theta_{t};X,Y(t))=Y(t)-\theta_{t}$ for $t\in\{0,1\}$.
We estimate $\theta_{1}$ and $\theta_{0}$ separately. For each $t\in\{0,1\}$, define $\delta_{i}=\mathds{1}(T_{i}=t)$, $O_{i}=X_{i}$, $M_{i}=Y_{i}(t)$, and $\pi(X_{i};\phi)=P(T_{i}=t\mid X_{i})$. The calibration function is $b(\theta_{t};X_{i})=U_{t}(\theta_{t};X_{i},\hat{Y}_{i}(t))=\hat{Y}_{i}(t)-\theta_{t}$, where $\hat{Y}_{i}(t)$ are cross-fitted predictions.
For the treated group ($t=1$), the calibration weights $\{\omega_{1i}\}_{i=1}^{N}$ are obtained by solving
$\min_{\omega_{11},\ldots,\omega_{1N}}\sum_{i=1}^{N}T_{i}\,G(\omega_{1i}),$ (6.1)
$\text{subject to}\quad\sum_{i=1}^{N}T_{i}\,\omega_{1i}=N,$ (6.2)
[…]
6.2 Semi-supervised learning
In many applied settings, the outcome variable is expensive or difficult to obtain while covariates are readily available. For example, in electronic health records research, clinical diagnoses require expert review but demographic and laboratory covariates are routinely recorded (Gronsbell and Cai, 2018; Chakrabortty and Cai, 2018). Semi-supervised methods aim to exploit the large pool of unlabeled covariates to improve efficiency over supervised analysis that uses only the labeled sample.
Formally, a labeled dataset $\mathcal{L}=\{(x_{i},y_{i});i=1,\ldots,n\}$ with fully observed outcomes is supplemented by an unlabeled dataset $\mathcal{U}=\{x_{i};i=n+1,\ldots,n+N\}$. Here $Z_{i}=(X_{i},Y_{i})$, $O_{i}=X_{i}$, $M_{i}=Y_{i}$, and $\delta_{i}=1$ if $Y_{i}$ is observed. The target parameter $\theta$ solves $\mathbb{E}\{U(\theta;X,Y)\}=0$, and the calibration function is $b(\theta;X_{i})=U(\theta;X_{i},\hat{Y}_{i})$, where $\hat{Y}_{i}$ are cross-fitted predictions of $Y_{i}$ based on $X_{i}$.
The selection mechanism may be either MAR, with $P(\delta_{i}=1\mid X_{i})=\pi(X_{i};\phi)$ estimated by logistic regression, or MCAR, with $P(\delta_{i}=1)=n/(n+N)$. In both cases, the calibration weights are obtained from (5.2)–(5.4) and the estimator from Algorithm 1. Concretely, the proposed estimator of $\theta$ solves the weighted estimating equation
$\frac{1}{n+N}\sum_{i=1}^{n+N}\delta_{i}\,\omega_{i}\,U(\theta;x_{i},y_{i})=0,$ (6.5)
[…]
6.3 Missing covariates in regression
Suppose $(Y_{i},X_{i})$ with $X_{i}=(X_{1i},X_{2i})$ are related by a regression model $Y_{i}=\mu(X_{1i},X_{2i};\beta)+\epsilon_{i}$ with $\mathbb{E}(\epsilon_{i}\mid X_{i})=0$, and $X_{2i}$ is subject to missingness. Here $O_{i}=(X_{1i},Y_{i})$, $M_{i}=X_{2i}$, and $\delta_{i}=\mathds{1}(X_{2i}\text{ observed})$. Under MAR,
$\delta_{i}\perp X_{2i}\mid(X_{1i},Y_{i}),$
so that the probability of observing $X_{2i}$ may depend on both the fully observed covariates $X_{1i}$ and the response $Y_{i}$, but not on the missing covariate $X_{2i}$ itself. This conditioning on $Y_{i}$ distinguishes the missing-covariate setting from the missing-outcome case: the propensity score model is
$P(\delta_{i}=1\mid X_{1i},Y_{i})=\pi(X_{1i},Y_{i};\,\phi),$
which can be estimated by maximum likelihood, e.g.,
$\hat{\phi}=\arg\max_{\phi}\sum_{i=1}^{N}\bigl[\delta_{i}\log\{\pi(X_{1i},Y_{i};\phi)\}+(1-\delta_{i})\log\{1-\pi(X_{1i},Y_{i};\phi)\}\bigr].$
The estimating function is
$U(\beta;\,X_{1i},X_{2i},Y_{i})=\{Y_{i}-\mu(X_{1i},X_{2i};\beta)\}\,\frac{\partial}{\partial\beta}\mu(X_{1i},X_{2i};\beta),$
and the calibration function is $b(\beta;X_{1i},Y_{i})=U(\beta;X_{1i},\hat{X}_{2i},Y_{i})$, where $\hat{X}_{2i}$ are cross-fitted predictions of $X_{2i}$ based on $(X_{1i},Y_{i})$. The calibration weights and estimator of $\beta$ are obtained by applying (5.2)–(5.4) and Algorithm 1.
7 Simulation Study
We evaluate the proposed Hellinger distance (HD) and exponential tilting (ET) entropy calibration estimators across the three application settings. All simulations use $K=4$ folds for cross-fitting unless otherwise noted.
7.1 Causal inference
Design.
We consider a $2\times 2$ factorial design with two outcome regression models (OR1, OR2) and two propensity score models (PS1, PS2), using $M=500$ Monte Carlo replicates with sample size N=1, [^0] N=1{,}000. Covariates are generated as $\bm{x}_{i}=(x_{i1},x_{i2},x_{i3},x_{i4})^{\top}\sim N(0_{4},I_{4})$, and potential outcomes as $y_{i}(t)=m_{t}(\bm{x}_{i})+\varepsilon_{it}$, $\varepsilon_{it}\sim N(0,1)$, $t\in\{0,1\}$.
  • OR1 (linear): $m_{1}(x)=1+\mathbf{1}^{\top}\bm{x}$, $m_{0}(x)=\mathbf{1}^{\top}\bm{x},\mathbf{1}=(1,1,1,1)^{\top}$; true ATE =1=1.
  • OR2 (nonlinear): Let $h(x)=(x-1)^{3}-x^{2}+x/\{1+\exp(\mathrm{clip}(x))\}+10$ with $\mathrm{clip}(x)=\max\{\min(x,3),-3\}$ applied componentwise. Set $m_{1}(x)=10+\mathbf{1}^{\top}\bm{x}+\tfrac{1}{2}\mathbf{1}^{\top}h(\bm{x})$, $m_{0}(\bm{x})=\mathbf{1}^{\top}\bm{x}+\tfrac{1}{2}\mathbf{1}^{\top}h(\bm{x})$; true ATE =10=10.
  • PS1 (logistic linear): $\pi(\bm{x})=\mathrm{expit}(-0.25+x_{1}+0.5x_{2}-0.5x_{3}-0.1x_{4})$.
  • PS2 (logistic nonlinear): $\pi(\bm{x})=\mathrm{expit}(x_{1}-0.5x_{1}x_{2}-x_{3}^{2}+0.5x_{4}^{3})$.
The treatment indicator is $T_{i}\sim\mathrm{Bernoulli}\{\pi(x_{i})\}$ and the observed outcome is $Y_{i}=T_{i}Y_{i}(1)+(1-T_{i})Y_{i}(0)$. We compare HD and ET with seven existing methods: inverse probability weighting (IPW; Horvitz and Thompson, 1952), entropy balancing (EBPS; Hainmueller, 2012), optimal covariate balancing (oCBPS; Fan et al., 2022), covariate balancing propensity score (CBPS; Imai and Ratkovic, 2014), empirical balancing calibration weighting (EBCW; Chan et al., 2016), and AIPW with either a linear model or GAM for the outcome regression, denoted AIPW (LM) and AIPW (GAM).
Figure 1: Estimation of Average Treatment Effects (ATE) under four scenarios: under OR1(2), the outcome regression (OR) model is correctly specified (misspecified); under PS1(2), the propensity score model is correctly specified (misspecified). The horizontal red line represents the true ATE.
Results.
Figure 1 summarizes the Monte Carlo distributions across four specification regimes. Under OR1PS1 (both models correct), all methods show negligible bias and similar variability, consistent with the local efficiency result in Remark 1: when both models are correctly specified, the proposed estimator matches AIPW.
Under OR1PS2 (PS misspecified, OR correct), IPW and oCBPS exhibit substantial bias because they rely exclusively on the PS model, whereas HD, ET, and the remaining calibration-based and augmented estimators maintain low bias, confirming the OR-based consistency pathway in Corollary 2.
The most informative comparison is OR2PS1 (OR misspecified, PS correct), which isolates the variance-reduction mechanism identified in Corollary 3. Here, HD and ET display noticeably smaller interquartile ranges than all competitors, including both AIPW variants. Among the AIPW estimators, AIPW (GAM) shows a clear efficiency gain over AIPW (LM), as the more flexible outcome model partially compensates for misspecification; nevertheless, both AIPW versions remain less efficient than HD and ET. The weighting-only methods (IPW, EBPS, oCBPS, CBPS, EBCW) show substantially wider boxplots, reflecting their lack of outcome-model augmentation.
Under joint misspecification (OR2PS2), all methods are potentially inconsistent. Nevertheless, HD and ET remain the most robust, exhibiting the smallest bias and most concentrated sampling distributions, which suggests that the entropy calibration framework degrades gracefully when neither model is fully correct. This is somewhat related to the global robustness of the HD discussed by Antoine and Dovonon (2021).
7.2 Semi-supervised learning
Design.
The target is the regression coefficient $\beta$ in $y_{i}=x_{i}^{\top}\beta+\epsilon_{i}$. Covariates are $x_{i}\sim N(0_{p},I_{p})$ with $p=4$, total sample size n+N=2,000n+N=2{,}000, and M=1, [^0] M=1{,}000 replications. The outcome is generated under two specifications:
  • OR1 (linear): $y_{i}=\alpha_{0}+\alpha_{1}^{\top}x_{i}+\epsilon_{i}$;
  • OR2 (nonlinear): $y_{i}=\alpha_{0}+\alpha_{1}^{\top}x_{i}+\alpha_{2}^{\top}\{x_{i}^{3}-x_{i}^{2}+\exp(x_{i})\}+\epsilon_{i}$,
where $(\alpha_{0},\alpha_{1}^{\top},\alpha_{2}^{\top})^{\top}=(1,\mathbf{1}^{\top},\mathbf{1}^{\top})^{\top}$ and $\epsilon_{i}\sim N(0,1)$. The true value of the target parameter $\beta$ is computed by generating data of size $10^{7}$. Note that under OR2 the data-generating model is nonlinear, so the linear working model $y_{i}=x_{i}^{\top}\beta+\epsilon_{i}$ is misspecified; nevertheless, $\beta$ remains a well-defined projection parameter that our method targets. Two selection mechanisms are considered: MAR, with $\mathrm{logit}(p_{i})=-1-x_{1i}-0.5x_{2i}+0.5x_{3i}+0.1x_{4i}$; and MCAR, with $p_{i}=n/(n+N)$. We compare HD and ET with the supervised estimator (Sup) and four semi-supervised methods: the projection-based estimator (PSSE; Song et al., 2024), density-ratio estimation (DRESS; Kawakita and Kanamori, 2013), efficient adaptive estimation (EASE; Chakrabortty and Cai, 2018), and the partial-information estimator (PI; Azriel et al., 2022).
Results.
Under OR model misspecification (OR2) with MAR (Figure 2), HD and ET achieve the highest efficiency among all methods for all five regression coefficients $\beta_{0},\ldots,\beta_{4}$. The improvement is most visible for $\beta_{0}$: the interquartile range of HD and ET is roughly half that of the supervised estimator, and noticeably smaller than those of PI, EASE, DRESS, and PSSE. The existing semi-supervised methods (PI, EASE, DRESS, PSSE) provide only modest gains over the supervised estimator under OR misspecification, because their efficiency improvements rely on the outcome model being approximately correct. In contrast, our proposed estimators benefit from the debiasing constraint, which compensates for outcome-model misspecification through the propensity-score information.
Under OR2 with MCAR (Figure 3), the same pattern holds: HD and ET dominate in efficiency. The MCAR mechanism simplifies the selection model, so the debiasing constraint is particularly effective, and the efficiency gains of HD and ET over competitors are even more pronounced.
Under a correctly specified OR model (OR1) with MCAR (Figure 4), all semi-supervised estimators—including HD, ET, PI, EASE, DRESS, and PSSE—are consistent and nearly as efficient as the supervised estimator, confirming that the proposed method does not sacrifice efficiency when the outcome model is correct. This is consistent with the local efficiency result (Remark 1): when both models are correctly specified, the calibration weights are close to unity and the estimator behaves like the full-sample supervised estimator.
Figure 2: Boxplots of the estimated linear regression coefficients $\bm{\beta}$ under OR2 (OR model misspecification) with MAR missingness.
Figure 3: Boxplots of the estimated linear regression coefficients $\bm{\beta}$ under OR2 (OR model misspecification) with MCAR missingness.
Figure 4: Estimation of the linear regression coefficients $\bm{\beta}$ under OR1 (when the OR model is correctly specified) and MCAR missingness mechanism.
7.3 Missing covariates in regression
Design.
We estimate $\beta=(\beta_{0},\beta_{1},\beta_{2})$ in the linear model $y_{i}=\beta_{0}+\beta_{1}x_{1i}+\beta_{2}x_{2i}+\epsilon_{i}$, where $x_{1i}\sim N(0,1)$, $x_{2i}\sim\mathrm{Bernoulli}(0.5)$, $\epsilon_{i}\sim N(0,1)$, and $x_{2i}$ is subject to missingness. Sample size is $N=500$ with M=1, [^0] M=1{,}000 replications. Two PS models govern the missingness of $x_{2i}$: PS1 (MAR), $\mathrm{logit}(\pi_{i})=-1+0.5x_{1i}+0.5y_{i}$; PS2 (MCAR), $\mathrm{logit}(\pi_{i})=-1$. Two OR models generate $y_{i}$:
  • OR1: $y_{i}=1+x_{1i}+2x_{2i}+\epsilon_{1i}$, $\epsilon_{1i}\sim N(0,1)$;
  • OR2: $y_{i}=0.5+2\sin(\pi x_{1i})-1.5\cos(2\pi x_{1i})+0.25x_{1i}^{3}+x_{2i}+\epsilon_{2i}$, $\epsilon_{2i}\sim N(0,4)$.
Cross-fitted predictions $\hat{x}_{2i}$ are obtained via logistic regression, since $x_{2i}$ is binary. We compare the proposed HD estimator with four benchmarks: the full-sample estimator (Full, infeasible), the complete-case estimator (CC), the Horvitz–Thompson estimator (HT), and the AIPW estimator.
Results.
Table 2 reports bias, standard error, and RMSE ($\times 10$) across all scenarios. The full-sample estimator serves as a benchmark with negligible bias and smallest RMSE. The CC estimator shows substantial bias under MAR (OR1PS1 and OR2PS1), confirming that discarding incomplete cases introduces systematic error when the missingness depends on the observed data.
Under OR1PS1 (both models correct), HD achieves RMSE comparable to or slightly better than AIPW for all three coefficients (e.g., RMSE of $0.71$ vs. $0.75$ for $\beta_{0}$, and $1.12$ vs. $1.18$ for $\beta_{2}$). Both HT and AIPW reduce bias relative to CC but exhibit inflated standard errors.
The largest efficiency gain appears under OR2PS1 (OR misspecified, PS correct). Here, HD achieves RMSE of $1.31$, $0.92$, and $2.48$ for $\beta_{0}$, $\beta_{1}$, and $\beta_{2}$ respectively, compared with $1.93$, $1.44$, and $3.79$ for AIPW—a reduction of approximately $32\%$, $36\%$, and $35\%$. The HT estimator, which does not use outcome information, performs comparably to AIPW or worse, highlighting the value of incorporating the calibration function even when the outcome model is imperfect.
Under MCAR (PS2), all weighting methods perform comparably because the selection mechanism is simple and correctly specified by all methods, with HD showing a slight advantage in OR1PS2 (e.g., RMSE of $0.87$ vs. $0.90$ for $\beta_{2}$) and essentially equivalent performance in OR2PS2.
Model Method $\beta_{0}=1$ $\beta_{1}=1$ $\beta_{2}=2$
Bias SE RMSE Bias SE RMSE Bias SE RMSE
Full -0.01 0.43 0.43 -0.03 0.33 0.33 0.04 0.63 0.64
$C$ 3.83 1.00 3.96 -1.01 0.63 1.19 -0.94 1.16 1.50
OR1PS1 HT 0.14 1.38 1.39 -0.13 1.01 1.02 -0.09 1.79 1.79
AIPW -0.07 0.75 0.75 -0.03 0.82 0.82 0.16 1.17 1.18
HD -0.12 0.70 0.71 -0.08 0.74 0.75 0.35 1.07 1.12
Full -0.01 0.43 0.43 -0.03 0.33 0.33 0.04 0.63 0.64
$C$ 0.01 0.85 0.85 -0.04 0.62 0.62 0.05 1.24 1.24
OR1PS2 HT 0.00 0.77 0.77 -0.04 0.62 0.62 0.05 1.24 1.24
AIPW -0.02 0.64 0.64 -0.03 0.50 0.50 0.09 0.90 0.90
HD -0.04 0.63 0.63 -0.03 0.50 0.50 0.13 0.86 0.87
Full 0.00 0.94 0.94 0.00 0.85 0.85 0.01 1.30 1.30
$C$ 11.44 1.56 11.54 -2.66 1.44 3.02 -2.05 2.00 2.86
OR2PS1 HT 0.51 1.88 1.95 -0.98 2.33 2.53 -0.15 2.77 2.77
AIPW -0.03 1.93 1.93 -0.09 1.43 1.44 0.21 3.78 3.79
HD -0.09 1.31 1.31 0.00 0.92 0.92 0.35 2.45 2.48
Full -0.02 1.22 1.22 -0.03 1.02 1.02 0.04 1.72 1.72
$C$ 0.00 2.20 2.20 -0.13 2.03 2.04 0.05 3.30 3.30
OR2PS2 HT -0.01 1.81 1.81 -0.14 2.02 2.03 0.05 3.32 3.32
AIPW -0.02 1.78 1.78 -0.02 1.07 1.07 0.08 3.26 3.26
HD -0.08 1.84 1.84 -0.02 1.08 1.08 0.20 3.36 3.37
Table 2: Bias ($\times 10$), standard error (SE) ($\times 10$) and root mean square error (RMSE) ($\times 10$) for estimators under missing covariates in regression model setting.
8 Real-Data Applications
8.1 Causal inference: the LaLonde job-training study
We assess the proposed estimators by estimating the treatment effect of the National Supported Work (NSW) Demonstration, a randomized job-training program conducted in the mid-1970s (LaLonde, 1986; Dehejia and Wahba, 1999). Following the standard benchmark design, we pool the NSW experimental sample with two large observational control groups—the Panel Study of Income Dynamics (PSID) and the Current Population Survey (CPS)—into a single sample. Units are classified by a group indicator $G_{i}\in\{1,2,3,4\}$: NSW treated ($N_{1}=185$), NSW controls ($N_{2}=260$), PSID ($N_{3}=2{,}490$), and CPS ($N_{4}=15{,}992$).
The target estimand is the average treatment effect for the NSW experimental population,
$\mathrm{ATE}_{\mathrm{NSW}}\equiv\mathbb{E}\{Y(1)-Y(0)\mid\text{NSW participants}\}.$
We estimate generalized propensity scores $\pi_{ig}=\Pr(G_{i}=g\mid x_{i})$ via multinomial logistic regression and define tilting weights $r_{ig}=\pi_{i,\mathrm{NSW}}/\pi_{ig}$, where $\pi_{i,\mathrm{NSW}}=\pi_{i1}+\pi_{i2}$. The covariates used for propensity score estimation are age, years of education, indicators for Black and Hispanic ethnicity, marital status, an indicator for no high school degree, and real earnings in 1974 and 1975 (re74 and re75). These are the standard covariates used in the LaLonde benchmark literature (Dehejia and Wahba, 1999). Cross-fitted outcome predictions $\hat{y}_{i}(g)$ are obtained using GAM with $K=4$ folds. For each candidate control group $g\in\{2,3,4\}$, calibration weights are constructed by solving (5.2)–(5.4) with the two-loop procedure, and the ATE is estimated as $\widehat{\mathrm{ATE}}_{g}=\bar{Y}_{1}-\hat{\theta}_{g}$, where $\bar{Y}_{1}$ is the mean outcome among NSW treated units.
Results.
Table 3 reports point estimates, standard errors, and evaluation bias (EB), defined as the difference from the NSW experimental benchmark $\hat{\theta}_{\mathrm{NSW}}=1794$. Using the NSW experimental data, all estimators yield estimates close to the benchmark, as expected under randomization. Using PSID or CPS as external controls, the unweighted difference in means produces large EB (17,000-17{,}000 and 10,292-10{,}292, respectively), reflecting substantial covariate imbalance between the NSW treated group and the observational control samples. Among all weighting methods, the proposed ET estimator most closely reproduces the experimental benchmark in both external-control settings, achieving the smallest absolute EB (139-139 for PSID; 179-179 for CPS) with standard errors comparable to competing methods. For the PSID comparison, the next-best method is IPW with EB =320=-320, more than twice the absolute bias of ET. For the CPS comparison, AIPW (GAM) achieves EB =111=-111, which is slightly smaller in absolute value than ET’s 179-179, but ET has a smaller standard error ($635$ vs. $656$), so the two are broadly comparable. Overall, ET provides the most consistently accurate reproduction of the experimental benchmark across both external-control settings.
Figure 5 compares weighted covariate distributions (age, education, re74, and re75) for PSID and CPS controls against the empirical distribution of the combined NSW sample. ET achieves the closest overall covariate balance across all four variables in both panels. For the PSID controls, IPW shows noticeable residual imbalance in age and re74, with the weighted CDF deviating visibly from the NSW target, while CBPS and EBCW improve upon IPW but still exhibit some discrepancy in the tails of the earnings distributions. For the CPS controls, the covariate distributions are initially closer to the NSW sample, so all weighting methods perform reasonably well, though ET and EBCW achieve the tightest alignment. These patterns are consistent with the numerical results in Table 3: methods that achieve better covariate balance tend to produce smaller evaluation bias.
(a) Treatment effects
NSW Data PSID Data CPS Data
Estimators Estimates SE Estimates EB SE Estimates EB SE
Unweighted 1794 671 -15205 -17000 657 -8498 -10292 583
IPW 1796 673 1474 -320 901 1064 -730 644
CBPS 1636 687 1389 -405 886 1288 -506 641
EBCW 1792 666 2316 522 799 1284 -510 633
AIPW (LM) 1779 673 1283 -511 909 1291 -503 652
AIPW (GAM) 1780 672 1426 -368 893 1683 -111 656
ET 1787 670 1655 -139 912 1615 -179 635
Table 3: Comparison of Average Treatment Effect Estimators on the NSW Experimental and PSID and CPS observational data. Notes: SE= standard error, EB = Evaluation Bias.
Figure 5: Weighted covariate distributions for the LaLonde (1986) data.
8.2 Semi-supervised learning: NHANES fasting glucose
We use the 2017–2018 National Health and Nutrition Examination Survey (NHANES) to estimate the association between fasting plasma glucose and cardiometabolic risk factors. NHANES is a nationally representative health survey conducted by the U.S. Centers for Disease Control and Prevention that collects detailed demographic, anthropometric, and clinical information on participants. Fasting glucose is measured only for participants attending the morning fasting examination, creating a natural semi-supervised setting: covariates are available for the full sample of N=6,230N=6{,}230 individuals, while glucose is observed for n=2,533n=2{,}533 (the labeled sample).
The covariates included in the regression model are age (years), sex (female indicator), race/ethnicity (indicators for Hispanic, non-Hispanic White, non-Hispanic Black, Asian, and Other), body mass index (BMI, kg/m2), systolic blood pressure (SBP, mmHg), and diastolic blood pressure (DBP, mmHg). These variables are standard cardiometabolic risk factors and are fully observed for all participants.
To better reflect practical semi-supervised scenarios with limited labeled data, we randomly subsample 50% of the labeled set, yielding n=1,266n=1{,}266 labeled and Nn=3,697N-n=3{,}697 unlabeled observations ($\approx$74% unlabeled). We compare the proposed ET estimator with the supervised (OLS) estimator, the density-ratio estimator (DRESS; Kawakita and Kanamori, 2013), and the projection-based estimator (PSSE; Song et al., 2024), using $K=3$ folds for cross-fitting.
Results.
Table 4 reports point estimates, standard errors, confidence interval widths (CIW), and asymptotic relative efficiency (ARE) relative to the supervised estimator. ARE values greater than one indicate that the semi-supervised estimator is more efficient (has smaller asymptotic variance) than the supervised baseline.
Method Inter. age sexF rHisp rWhite rBlack rAsian rOther BMI SBP DBP
Supervised Est 72.31 0.43 -5.86 -0.79 -6.33 -9.21 -1.41 -6.71 0.82 0.10 -0.12
SE 7.94 0.06 1.99 4.16 3.06 3.26 3.74 4.82 0.14 0.07 0.08
CIW 31.13 0.23 7.80 16.32 12.01 12.78 14.66 18.88 0.55 0.26 0.32
DRESS Est 71.55 0.39 -4.24 -2.16 -6.33 -7.90 -2.51 -6.56 0.64 0.16 -0.12
SE 10.86 0.07 1.98 5.28 3.61 3.64 3.83 3.77 0.17 0.11 0.06
CIW 42.57 0.27 7.75 20.69 14.16 14.28 15.02 14.79 0.67 0.42 0.22
ARE 0.54 0.70 1.01 0.62 0.72 0.80 0.95 1.63 0.65 0.36 2.17
PSSE Est 71.98 0.40 -4.55 -2.16 -6.91 -8.25 -2.73 -7.22 0.66 0.16 -0.14
SE 10.80 0.07 1.97 5.25 3.60 3.63 3.82 3.76 0.17 0.11 0.06
CIW 42.34 0.27 7.72 20.56 14.11 14.24 14.97 14.75 0.67 0.42 0.22
ARE 0.54 0.71 1.02 0.63 0.73 0.81 0.96 1.64 0.66 0.37 2.19
ET Est 77.08 0.40 -4.97 -0.03 -5.21 -7.55 -1.70 -5.21 0.67 0.12 -0.15
SE 8.07 0.05 1.59 3.83 2.69 2.86 3.07 2.96 0.14 0.08 0.08
CIW 31.63 0.20 6.24 15.02 10.55 11.22 12.05 11.59 0.53 0.31 0.30
ARE 0.97 1.33 1.56 1.18 1.29 1.30 1.48 2.65 1.04 0.70 1.16
Table 4: Comparison of semi-supervised estimators using a 50% labeled subsample. Notes: Est = estimates, SE = standard error. CIW = confidence interval width, Proportion=50%, ARE= estimated asymptotic relative efficiency relative to the supervised estimator, we use K=3 folds for cross-fitting.
The proposed ET estimator achieves the largest efficiency gains across methods. For the age coefficient, ET yields a standard error of $0.05$ compared with $0.06$ for the supervised estimator (ARE =1.33=1.33), while DRESS and PSSE both have SE =0.07=0.07 (ARE $\approx 0.70$), indicating that these competing methods are actually less efficient than the supervised estimator for this coefficient. A similar pattern holds for the sex coefficient: ET achieves SE =1.59=1.59 versus the supervised SE =1.99=1.99 (ARE =1.56=1.56), whereas DRESS and PSSE have SE $\approx 1.98$ with ARE barely above $1.0$. The largest efficiency gain for ET appears in the “Other race” coefficient, where ARE =2.65=2.65 corresponds to a confidence interval width of $11.59$ compared with $18.88$ for the supervised estimator—a reduction of nearly 39%.
DRESS and PSSE produce ARE values below one for several coefficients (e.g., intercept, Hispanic, SBP), meaning they are less efficient than the supervised estimator. This occurs because these methods rely on the outcome prediction model, and when the prediction is imperfect, the additional variability from incorporating unlabeled data can outweigh the efficiency gain. In contrast, ET’s debiasing constraint provides a safeguard against this phenomenon: by incorporating the propensity-score information, it ensures that the unlabeled data contribute to efficiency even when the outcome model is imperfect, consistent with the doubly robust property established in Section 4.
9 Conclusion
We have proposed a unified calibration-weighting framework for estimating parameters defined by general estimating equations with partially observed data. The method constructs weights by minimizing a generalized entropy subject to balancing constraints that incorporate cross-fitted machine-learning predictions and debiasing constraints based on a working propensity-score model. The resulting estimator is doubly robust and attains the semiparametric efficiency bound when both the outcome-regression and propensity-score models are correctly specified. A distinctive theoretical property, established in Corollary 3, is that the proposed estimator has smaller asymptotic variance than the classical AIPW estimator whenever the OR model is misspecified, because the debiasing constraint enriches the calibration space and reduces the residual variance of the estimating function after projection.
This theoretical advantage is confirmed empirically. In the causal inference simulations, HD and ET exhibit noticeably smaller variability than all competitors under outcome-model misspecification (OR2PS1), while matching AIPW under correct specification. In the missing-covariates setting, the proposed HD estimator reduces RMSE by approximately 32–36% relative to AIPW when the OR model is misspecified and the PS model is correct. In the NHANES semi-supervised application, the ET estimator achieves asymptotic relative efficiency gains of up to 2.65 over the supervised estimator, whereas competing semi-supervised methods sometimes perform worse than the supervised baseline. These results collectively demonstrate that generalized entropy calibration provides a practical and reliable approach to inference with partially observed data, particularly in settings where the outcome model may be misspecified.
Several other extensions merit investigation. First, the current framework assumes a missing-at-random mechanism; extending the approach to handle non-ignorable missingness, where the probability of observation depends on the unobserved values themselves, would substantially broaden its applicability. Second, when the dimension of the calibration function $b(O)$ is large relative to the sample size, the entropy calibration problem may become ill-conditioned; incorporating regularization into the dual objective or employing dimension-reduction techniques for the calibration constraints could address this challenge. More broadly, the proposed framework provides a general tool for inference with partially observed data and can be adapted to problems involving data integration across multiple sources.
Appendix
In this supplementary material, we provide all regularity conditions (Section A), proofs of the theoretical results stated in the main paper (Section B), and asymptotic theory for cross-fitted GEC Estimators (Section C).
Appendix A Assumptions
We organize the assumptions into three groups: conditions for the consistency of the nuisance parameters $\hat{\varphi}$ and $\hat{\lambda}$ (Assumptions 12), conditions for the linearization of the weighted estimating equation (Assumptions 34), and conditions for the joint asymptotic normality of $\hat{\theta}_{\omega}$ (Assumptions 57).
Consistency of $\hat{\varphi}$ and $\hat{\lambda}$.
Assumption 1 (PS model regularity).
The limiting function $\mathbb{E}\{\nabla\ell(\varphi)\}$ is continuously differentiable with respect to $\varphi$ on a compact set $\mathcal{G}_{\varphi}$ containing $\varphi^{*}$.
Assumption 2 (Calibration regularity).
  1. (i)
    $b_{i}$ and $O_{i}$ have compact support.
  2. (ii)
    $\mathbb{E}(\delta\,s_{i}^{*\top}s_{i}^{*})$ is nonsingular, where $s_{i}^{*}=s_{i}(\varphi^{*})$.
  3. (iii)
    The convex conjugate $F$ of $G$ is strictly convex.
Linearization.
Assumption 3 (Smoothness of limiting objectives).
The limiting functions $\mathbb{E}\{\nabla\rho_{G}(\lambda)\}$ and $\mathbb{E}\{\nabla\ell(\varphi)\}$ are differentiable, and their Hessians $\mathbb{E}\{\nabla^{2}\rho_{G}(\lambda)\}$ and $\mathbb{E}\{\nabla^{2}\ell(\varphi)\}$ are continuous on compact sets $\mathcal{G}_{\lambda}\ni\lambda^{*}$ and $\mathcal{G}_{\varphi}\ni\varphi^{*}$, respectively.
Assumption 4 (Non-degeneracy).
The matrices $\mathbb{E}\{\nabla^{2}\ell(\varphi^{*})\}$ and $\mathbb{E}\{\nabla^{2}\rho_{G}(\lambda^{*})\}$ are non-singular.
Joint asymptotic normality.
Assumption 5 (Existence and identifiability).
Let $\alpha=(\theta^{\top},\lambda^{\top},\varphi^{\top})^{\top}\in\mathbb{R}^{2q+r+1}$ and
$\hat{Q}(\alpha)\;=\;\bigl(\nabla\ell(\varphi)^{\top},\;\nabla\rho_{G}(\lambda)^{\top},\;\hat{U}^{\omega}(\theta)^{\top}\bigr)^{\top}.$
The equation $\hat{Q}(\alpha)=0$ has a unique solution $\hat{\alpha}=(\hat{\theta}^{\top},\hat{\lambda}^{\top},\hat{\varphi}^{\top})^{\top}$, and there exists a function $\mathbb{E}\{\hat{Q}(\alpha)\}$ such that $\hat{Q}(\alpha)\to\mathbb{E}\{\hat{Q}(\alpha)\}$ uniformly as $N\to\infty$ and $\mathbb{E}\{\hat{Q}(\alpha)\}=0$ has a unique solution $\alpha^{*}=(\theta^{*\top},\lambda^{*\top},\varphi^{*\top})^{\top}$.
Assumption 6 (Smoothness of the joint system).
The limiting function $\mathbb{E}\{\hat{Q}(\alpha)\}$ is differentiable and $\mathbb{E}\{\partial\hat{Q}(\alpha)/\partial\alpha\}$ is continuous on a compact set $\mathcal{G}=\mathcal{G}_{\theta}\times\mathcal{G}_{\lambda}\times\mathcal{G}_{\varphi}$ containing $\alpha^{*}$.
Assumption 7 (Non-singularity of the joint Jacobian).
The matrix $\mathbb{E}\{\partial\hat{Q}(\alpha)/\partial\alpha\}\big|_{\alpha=\alpha^{*}}$ is non-singular.
Appendix B Proofs
B.1 Proof of Lemma S.1 (Consistency of $\hat{\varphi}$ and $\hat{\lambda}$)
Lemma B.1 (Consistency).
Under Assumptions 12, $\hat{\varphi}\xrightarrow{p}\varphi^{*}$ and $\hat{\lambda}\xrightarrow{p}\lambda^{*}$, where $\varphi^{*}=\arg\max_{\varphi}E\{\ell(\varphi)\}$ and $\lambda^{*}=\arg\min_{\lambda}E\{\hat{\rho}_{G}(\lambda)\}$.
Proof.
Step 1: Consistency of $\hat{\varphi}$.  The pseudo-true value $\varphi^{*}$ minimizes the Kullback–Leibler divergence
$\mathrm{KL}(\varphi)\;=\;\mathbb{E}\!\left[\log\frac{g(\delta\mid O)}{f(\delta\mid O;\,\varphi)}\right],$
where $g(\delta\mid O)$ is the true conditional density of $\delta$ given $O$ and $f(\delta\mid O;\,\varphi)=\pi(O;\varphi)^{\delta}\{1-\pi(O;\varphi)\}^{1-\delta}$ is the working model. By Assumption 1 and Theorem 2.2 of White (1982), $\hat{\varphi}\xrightarrow{p}\varphi^{*}$.
Step 2: Uniform convergence of $\nabla\hat{\rho}_{G}$.  By the consistency of $\hat{\varphi}$ and the delta method,
$\nabla\hat{\rho}_{G}(\lambda)\;=\;\nabla\rho_{G}(\lambda)+o_{p}(1)$ (B.1)
[…]
$\sup_{\lambda\in\mathcal{G}_{\lambda}}\|\nabla\hat{\rho}_{G}(\lambda)-\nabla\rho_{G}(\lambda)\|\;=\;o_{p}(1).$ (B.2)
Step 3: Uniqueness of $\lambda^{*}$ and consistency of $\hat{\lambda}$.  Suppose $\nabla\rho_{G}(\lambda_{1})=\nabla\rho_{G}(\lambda_{2})=0$. Then
$\mathbb{E}\!\bigl[\delta\,\{g^{-1}(\lambda_{1}^{\top}s^{*})-g^{-1}(\lambda_{2}^{\top}s^{*})\}(\lambda_{1}^{\top}s^{*}-\lambda_{2}^{\top}s^{*})\bigr]=0.$
Since $g^{-1}$ is strictly increasing, the integrand is non-negative, so $\lambda_{1}^{\top}s^{*}=\lambda_{2}^{\top}s^{*}$ a.s. for $\delta=1$. Non-singularity of $\mathbb{E}(\delta\,s^{*\top}s^{*})$ (Assumption 2(ii)) implies $\lambda_{1}=\lambda_{2}$. Combined with (B.2), Theorem 5.9 of Van der Vaart (2000) gives $\hat{\lambda}\xrightarrow{p}\lambda^{*}$. ∎
B.2 Proof of Theorem 4.1 (Linearization)
Proof.
[…]
Step 1: $\sqrt{N}$-rate for $\hat{\varphi}$ and $\hat{\lambda}$.  By a mean-value expansion of $\nabla\ell(\hat{\varphi})=0$ around $\varphi^{*}$ and Assumptions 34,
$\hat{\varphi}-\varphi^{*}\;=\;-\bigl[\mathbb{E}\{\nabla^{2}\ell(\varphi^{*})\}\bigr]^{-1}\,\nabla\ell(\varphi^{*})\;+\;o_{p}(N^{-1/2}),$ (B.3)
where $\nabla\ell(\varphi)=N^{-1}\sum_{i=1}^{N}(\delta_{i}/\pi_{i}-1)\,h_{i}(\varphi)$. An analogous expansion of $\nabla\hat{\rho}_{G}(\hat{\lambda})=0$ yields
$\hat{\lambda}-\lambda^{*}\;=\;-\bigl[\mathbb{E}\{\nabla^{2}\rho_{G}(\lambda^{*})\}\bigr]^{-1}\,\nabla\rho_{G}(\lambda^{*})\;+\;o_{p}(N^{-1/2}).$ (B.4)
Step 2: Expansion of $\hat{U}^{\omega}(\theta)$ around $\lambda^{*}$.  By the mean value theorem and (B.4),
$\hat{U}^{\omega}(\theta)$ $=\frac{1}{N}\sum_{i=1}^{N}\delta_{i}\,\omega_{i}^{*}(\lambda^{*})\,U(\theta;z_{i})\;+\;\left(\frac{1}{N}\sum_{i=1}^{N}\delta_{i}\,U(\theta;z_{i})\,\frac{\partial}{\partial\lambda}\omega_{i}^{*}(\tilde{\lambda})^{\top}\right)(\hat{\lambda}-\lambda^{*})$
$=\frac{1}{N}\sum_{i=1}^{N}\delta_{i}\,\omega_{i}^{*}(\lambda^{*})\,U(\theta;z_{i})\;-\;\gamma^{*}\,\nabla\rho_{G}(\lambda^{*})\;+\;o_{p}(N^{-1/2}),$ (B.5)
where the second equality uses Randles (1982) and
$\gamma^{*}\;=\;\mathbb{E}\!\bigl[\delta\,f^{\prime}(\lambda^{*\top}S)\,U(\theta;Z)\,S^{\top}\bigr]\;\bigl\{\mathbb{E}\!\bigl[\delta\,f^{\prime}(\lambda^{*\top}S)\,S\,S^{\top}\bigr]\bigr\}^{-1},$ (B.6)
with $S=S(O;\varphi^{*})$ and $f^{\prime}=(g^{-1})^{\prime}$.
Step 3: Assembling the linearization.  Substituting the definition of $\nabla\rho_{G}(\lambda^{*})$ into (B.5) and rearranging gives
$\hat{U}^{\omega}(\theta)\;=\;\frac{1}{N}\sum_{i=1}^{N}\Bigl[\gamma^{*}\,s_{i}(\hat{\varphi})\;+\;\delta_{i}\,\omega_{i}^{*}(\lambda^{*},\hat{\varphi})\bigl\{U(\theta;z_{i})-\gamma^{*}\,s_{i}(\hat{\varphi})\bigr\}\Bigr]+o_{p}(N^{-1/2}),$
which is the claimed form $\tilde{U}^{\omega}(\lambda^{*},\hat{\varphi})+o_{p}(N^{-1/2})$. No assumption on the correctness of either the PS or OR model was used. ∎
B.3 Proof of Lemma 4.2
Proof.
Set $\lambda^{*}=(0^{\top},1)^{\top}$. Under the correct PS model, $P(\delta=1\mid O)=\pi(O;\varphi_{0})$, so
$\mathbb{E}\{\nabla\rho_{G}(\lambda^{*})\}$ $=\mathbb{E}\!\left[\delta\,g^{-1}\!\bigl(g(\pi^{-1}(O;\varphi_{0}))\bigr)\,s^{*}\;-\;s^{*}\right]=\mathbb{E}\!\left[\frac{\delta}{\pi(O;\varphi_{0})}\,s^{*}-s^{*}\right]=0,$
where the last equality follows from $\mathbb{E}(\delta\mid O)=\pi(O;\varphi_{0})$. Since the solution is unique (Lemma B.1), $\lambda^{*}=(0^{\top},1)^{\top}$ is the probability limit, giving $\omega_{i}^{*}(\lambda^{*},\varphi_{0})=\pi^{-1}(O_{i};\varphi_{0})$. ∎
B.4 Proof of Corollary 1
Proof.
[…]
$\quad+o_{p}(N^{-1/2}),$ (B.7)
where $\bar{\varphi}$ lies between $\varphi^{*}$ and $\hat{\varphi}$. Substituting (B.3) and applying Randles (1982), the second term becomes $\kappa^{*}\,\nabla\ell(\varphi^{*})+o_{p}(N^{-1/2})$, where
$\kappa^{*}\;=\;\mathbb{E}\!\left[\frac{\partial}{\partial\varphi}\Bigl\{\gamma^{*}S(\varphi^{*})+\delta\,\omega^{*}(\lambda^{*},\varphi^{*})\bigl(U(\theta;Z)-\gamma^{*}S(\varphi^{*})\bigr)\Bigr\}\right]\!\left\{\mathbb{E}\!\left[\frac{1-\pi(O;\varphi^{*})}{\pi(O;\varphi^{*})}\,h(\varphi^{*})\,h^{\top}(\varphi^{*})\right]\right\}^{\!-1}\!.$ (B.8)
Step 2: Influence function.  Combining the above, we obtain the representation
$\hat{U}^{\omega}(\theta)\;=\;\frac{1}{N}\sum_{i=1}^{N}d_{i}\;+\;o_{p}(N^{-1/2}),$ (B.9)
where the influence function is
$d_{i}\;=\;\gamma^{*}s_{i}(\varphi^{*})\;+\;\delta_{i}\,\omega_{i}^{*}(\lambda^{*},\varphi^{*})\bigl\{U(\theta_{0};z_{i})-\gamma^{*}s_{i}(\varphi^{*})\bigr\}\;+\;\Bigl(1-\frac{\delta_{i}}{\pi(O_{i};\varphi^{*})}\Bigr)\kappa^{*}\,h_{i}(\varphi^{*}).$
Step 3: Asymptotic normality.  A standard mean-value expansion of $\hat{U}^{\omega}(\hat{\theta}_{\omega})=0$ around $\theta_{0}$ gives
$\hat{\theta}_{\omega}-\theta_{0}\;=\;-\tau_{1}^{-1}\;\frac{1}{N}\sum_{i=1}^{N}d_{i}\;+\;o_{p}(N^{-1/2}),$
where $\tau_{1}=\mathbb{E}\{\partial U(\theta_{0};Z)/\partial\theta\}$. Under the correct PS model, $\lambda^{*}=(0^{\top},1)^{\top}$ (Lemma 4.2), so $\omega_{i}^{*}(\lambda^{*},\varphi^{*})=\pi^{-1}(O_{i};\varphi_{0})$ and $\gamma^{*}$ simplifies to
$\gamma^{*}\;=\;\mathbb{E}\!\bigl[U(\theta_{0};Z)\,S^{\top}(\varphi^{*})\bigr]\;\bigl\{\mathbb{E}\!\bigl[S(\varphi^{*})\,S^{\top}(\varphi^{*})\bigr]\bigr\}^{-1}.$
Step 4: Variance calculation.  By the law of total variance,
$V_{1}=\mathrm{Var}(d_{i})=\mathrm{Var}\!\bigl\{\mathbb{E}(d_{i}\mid O_{i})\bigr\}+\mathbb{E}\!\bigl\{\mathrm{Var}(d_{i}\mid O_{i})\bigr\}.$
Since $\mathbb{E}(\delta_{i}\mid O_{i})=\pi(O_{i};\varphi_{0})$, the conditional mean is $\mathbb{E}(d_{i}\mid O_{i})=\mathbb{E}\{U(\theta_{0};Z)\mid O\}$ and the conditional variance is
$\mathrm{Var}(d_{i}\mid O_{i})\;=\;\left(\frac{1}{\pi(O_{i};\varphi_{0})}-1\right)\bigl\|U(\theta_{0};Z_{i})-\gamma^{*}S_{i}-\kappa^{*}h_{i}\bigr\|^{2},$
yielding the stated form of $V_{1}$. ∎
B.5 Proof of Lemma 4.3
Proof.
If $\mathbb{E}\{U(\theta_{0};Z)\mid O\}\in\mathrm{span}\{b(O),\,g(\pi^{-1}(O;\varphi))\}$, the weighted least-squares normal equation defining $\gamma^{*}$ in (B.6) is solved by $\gamma^{*}=(I_{q}\;\;0)$, so $\gamma^{*}S(O;\varphi^{*})=b^{*}(O)$.
To show $\kappa^{*}=0$, note that the numerator of (B.8) involves
$\mathbb{E}\!\left[\frac{\partial}{\partial\varphi}\Bigl\{b^{*}(O)+\delta\,\omega^{*}\!\bigl(U(\theta;Z)-b^{*}(O)\bigr)\Bigr\}\right].$
Conditioning on $(O,\delta)$ and using $\mathbb{E}\{U(\theta_{0};Z)\mid O\}=b^{*}(O)$, the inner expectation becomes
$\frac{\partial}{\partial\varphi}\Bigl\{b^{*}(O)+\delta\,\omega^{*}\!\bigl(b^{*}(O)-b^{*}(O)\bigr)\Bigr\}=0.$
Hence the numerator vanishes and $\kappa^{*}=0$. ∎
B.6 Proof of Corollary 2
Proof.
By Lemma 4.3, $\kappa^{*}=0$ and $\gamma^{*}S(O;\varphi^{*})=b^{*}(O)$, so the influence function (B.9) simplifies to
$d_{i}\;=\;b^{*}(O_{i})\;+\;\delta_{i}\,\omega_{i}^{*}(\lambda^{*},\varphi^{*})\bigl\{U(\theta_{0};z_{i})-b^{*}(O_{i})\bigr\}.$
Derivation of $\tau_{1}$.  Using the balancing constraint $\mathbb{E}\{\delta\,\omega^{*}\,b^{*}(O)\}=\mathbb{E}\{b^{*}(O)\}$,
$\tau_{1}=\mathbb{E}\!\left[\frac{\partial}{\partial\theta}\bigl\{\delta\,\omega^{*}\,U(\theta_{0};Z)\bigr\}\right]=\frac{\partial}{\partial\theta}\,\mathbb{E}\!\bigl[b^{*}(O)\bigr]=\mathbb{E}\!\left[\frac{\partial}{\partial\theta}\,U(\theta_{0};Z)\right].$
Derivation of $\bar{V}_{1}$.  By the law of total variance and the balancing constraint,
$\mathbb{E}(d_{i}\mid O_{i})$ $=b^{*}(O_{i})+\mathbb{E}(\delta_{i}\mid O_{i})\,\omega_{i}^{*}\bigl\{b^{*}(O_{i})-b^{*}(O_{i})\bigr\}=\mathbb{E}\{U(\theta_{0};Z)\mid O_{i}\},$
$\mathrm{Var}(d_{i}\mid O_{i})$ $=\delta_{i}\,\{\omega_{i}^{*}\}^{2}\,\mathrm{Var}\{U(\theta_{0};Z)\mid O_{i}\}.$
Therefore, $\bar{V}_{1}=\mathrm{Var}\!\bigl[\mathbb{E}\{U(\theta_{0};Z)\mid O\}\bigr]+\mathbb{E}\!\bigl[\delta\,\{\omega^{*}\}^{2}\,\mathrm{Var}\{U(\theta_{0};Z)\mid O\}\bigr]$. ∎
B.7 Proof of Corollary 3
Proof.
Define the weighted inner product $\langle f_{1},f_{2}\rangle_{w}=\mathbb{E}\!\bigl[\bigl(\pi^{-1}(O;\varphi_{0})-1\bigr)\,f_{1}\,f_{2}^{\top}\bigr]$ and the induced seminorm $\|f\|_{w}^{2}=\langle f,f\rangle_{w}$. The variance components differ only in the conditional second-moment term:
$V_{3}-V_{1}$ $=\bigl\|U(\theta_{0};Z)-b(O)\bigr\|_{w}^{2}-\bigl\|U(\theta_{0};Z)-\gamma^{*}S(O;\varphi_{0})-\kappa^{*}h(O;\varphi_{0})\bigr\|_{w}^{2}.$
Since $\mathrm{span}\{b(O)\}\subseteq\mathcal{S}=\mathrm{span}\{S(O;\varphi_{0}),\,h(O;\varphi_{0})\}$, the Pythagorean identity gives
$\bigl\|U-\Pi_{\{b\}}U\bigr\|_{w}^{2}\;=\;\bigl\|U-\Pi_{\mathcal{S}}U\bigr\|_{w}^{2}\;+\;\bigl\|\Pi_{\mathcal{S}}U-\Pi_{\{b\}}U\bigr\|_{w}^{2},$
where $\Pi_{\mathcal{A}}$ denotes $\|\cdot\|_{w}$-projection onto $\mathcal{A}$. The remainder term $\|\Pi_{\mathcal{S}}U-\Pi_{\{b\}}U\|_{w}^{2}\geq 0$ establishes $V_{1}\leq V_{3}$. Equality holds if and only if $\Pi_{\mathcal{S}}U=\Pi_{\{b\}}U$, which requires $g(\pi^{-1}(O;\varphi_{0}))$ and $h(O;\varphi_{0})\in\mathrm{span}\{b(O)\}$ a.s. ∎
Appendix C Theory for Cross-Fitted GEC Estimators
C.1 Overview and Motivation
The manuscript establishes the asymptotic theory of the generalized entropy calibration (GEC) estimator $\hat{\theta}_{\omega}$ in Theorem 4.1 and Corollaries 1–2 under the assumption that the calibration function $b(O_{i})$ is a known, fixed function. In practice, Section 5.2 replaces $b(O_{i})$ by a cross-fitted machine-learning prediction $\hat{b}_{i}^{(-k)}=U(\theta;O_{i},\hat{M}_{i}^{(-k)})$, and Remark 3 asserts that only mean-squared-error consistency is required for the cross-fitted predictor. In this section, we formalize the conditions under which the asymptotic results extend to the estimated calibration function.
We show that there is a fundamental asymmetry between the propensity-score (PS) and outcome-regression (OR) consistency pathways:
  • Under correct PS: only $L_{2}$-consistency of $\hat{b}$ is needed (no convergence rate).
  • Under correct OR (PS possibly wrong): the treatment depends on whether $\hat{b}$ is parametric or nonparametric.
    • Parametric $\hat{b}$: No rate condition is needed. The $O_{p}(N^{-1/2})$ estimation effect is absorbed into the influence function via a joint estimating equation argument, producing a correction term analogous to $\kappa^{*}$ in Corollary 1.
    • Nonparametric/ML $\hat{b}$: An explicit rate condition $\|\hat{b}-b^{*}\|_{2}=o_{p}(N^{-1/2})$ is required.
This asymmetry arises because the debiasing constraint drives the Lagrange multiplier $\lambda_{1}^{*}\to 0$ under a correct PS model, effectively decoupling the calibration weights from $b$ in the limit.
C.2 Setup and Notation
We adopt the notation of the manuscript. Let $b^{*}(O)=\mathbb{E}\{U(\theta_{0};Z)\mid O\}$ denote the optimal calibration function. With $K$-fold cross-fitting, we observe the cross-fitted calibration function $\hat{b}_{i}=\hat{b}_{i}^{(-k)}=U(\theta;O_{i},\hat{M}_{i}^{(-k)})$ for $i\in\mathcal{I}^{(k)}$, where $\hat{m}^{(-k)}$ is trained on the complement fold $\mathcal{S}^{(-k)}$.
Define:
$s_{i}^{*}$ $=\big(b^{*\top}(O_{i}),\;g(\pi^{-1}(O_{i};\varphi^{*}))\big)^{\top}\in\mathbb{R}^{q+1},$ (C.1)
$\hat{s}_{i}$ $=\big(\hat{b}_{i}^{\top},\;g(\hat{\pi}_{i}^{-1})\big)^{\top}\in\mathbb{R}^{q+1},$ (C.2)
$r_{i}$ $=\hat{b}_{i}-b^{*}(O_{i}).$ (C.3)
Let $\hat{\theta}_{\omega}^{\mathrm{cf}}$ denote the GEC estimator obtained from the cross-fitted calibration problem (5.2)–(5.4), and let $\hat{\theta}_{\omega}^{\mathrm{or}}$ denote the “oracle” GEC estimator using the true $b^{*}(O)$.
We impose the following cross-fitting condition.
Assumption 8 (Cross-fitting regularity).
$K$-fold cross-fitting is used with $K\geq 2$ fixed. For each fold $k=1,\ldots,K$, the predictor $\hat{m}^{(-k)}$ is trained on $\mathcal{S}^{(-k)}$ and is independent of the observations in fold $\mathcal{I}^{(k)}$.
C.3 Main Results
Proposition C.1 (Cross-fitted GEC under correct PS).
Under Assumptions 1–7, Assumption 8, and correct specification of the PS model $P(\delta=1\mid O)=\pi(O;\varphi_{0})$, suppose
$\mathbb{E}\big\|\hat{b}_{i}^{(-k)}-b^{*}(O_{i})\big\|^{2}=o_{p}(1),\quad k=1,\ldots,K.$ (C.4)
Then the cross-fitted estimator $\hat{\theta}_{\omega}^{\mathrm{cf}}$ satisfies
$\sqrt{N}(\hat{\theta}_{\omega}^{\mathrm{cf}}-\theta_{0})\xrightarrow{d}N\big(0,\;\tau_{1}^{-1}V_{1}(\tau_{1}^{-1})^{\top}\big),$
where $\tau_{1}$ and $V_{1}$ are defined as in Corollary 1. In particular, no convergence rate beyond $L_{2}$-consistency is required for $\hat{b}$.
Proposition C.2 (Cross-fitted GEC under correct OR: parametric calibration).
Under Assumptions 1–7, Assumption 8, and correct specification of the OR model in the sense that $\mathbb{E}\{U(\theta_{0};Z)\mid O\}\in\mathrm{span}\{b(O)\}$. Suppose the calibration function has a parametric form $b(O;\alpha)$ with $b^{*}(O)=b(O;\alpha_{0})$ for some finite-dimensional $\alpha_{0}\in\mathbb{R}^{p}$, and $\hat{\alpha}$ is a $\sqrt{N}$-consistent estimator satisfying
$\hat{\alpha}-\alpha_{0}=-A^{-1}\frac{1}{N}\sum_{i=1}^{N}\psi(z_{i},\delta_{i};\alpha_{0})+o_{p}(N^{-1/2}),$ (C.5)
where $A=\mathbb{E}\{\partial\psi/\partial\alpha\}$ is nonsingular. Then the cross-fitted estimator satisfies
$\sqrt{N}(\hat{\theta}_{\omega}^{\mathrm{cf}}-\theta_{0})\xrightarrow{d}N\big(0,\;\tau_{1}^{-1}V_{1}^{\mathrm{param}}(\tau_{1}^{-1})^{\top}\big),$
where
$V_{1}^{\mathrm{param}}=\mathrm{Var}\big(d_{i}^{\mathrm{param}}\big),$ (C.6)
with the modified influence function
$d_{i}^{\mathrm{param}}$ $=\gamma^{*}s_{i}(\varphi^{*})+\delta_{i}\omega_{i}^{*}(\lambda^{*},\varphi^{*})\big\{U(\theta_{0};z_{i})-\gamma^{*}s_{i}(\varphi^{*})\big\}$
$\quad+\big(1-\delta_{i}/\pi_{i}\big)\kappa^{*}h_{i}(\varphi^{*})-c_{\alpha}^{\top}A^{-1}\psi(z_{i},\delta_{i};\alpha_{0}),$ (C.7)
and
$c_{\alpha}=\mathbb{E}\big\{[\pi(O)\omega^{*}(O;\lambda^{*},\varphi^{*})-1]\,\dot{b}(O;\alpha_{0})\big\}\in\mathbb{R}^{p},$ (C.8)
with $\dot{b}(O;\alpha)=\partial b(O;\alpha)/\partial\alpha$.
If $\dot{b}(O;\alpha_{0})\in\mathrm{span}\{s^{*}(O)\}$, then $c_{\alpha}=0$ and the asymptotic variance reduces to $\tau_{1}^{-1}\bar{V}_{1}(\tau_{1}^{-1})^{\top}$ as in Corollary 2.
Proposition C.3 (Cross-fitted GEC under correct OR: nonparametric/ML calibration).
Under Assumptions 1–7, Assumption 8, and correct specification of the OR model in the sense that $\mathbb{E}\{U(\theta_{0};Z)\mid O\}\in\mathrm{span}\{b(O)\}$ where $b(O)$ is the population calibration function, suppose
$\big\|\hat{b}^{(-k)}-b^{*}\big\|_{2}=o_{p}(N^{-1/2}),\quad k=1,\ldots,K,$ (C.9)
where $\|\cdot\|_{2}$ denotes the $L_{2}(P)$-norm. Then the cross-fitted estimator satisfies
$\sqrt{N}(\hat{\theta}_{\omega}^{\mathrm{cf}}-\theta_{0})\xrightarrow{d}N\big(0,\;\tau_{1}^{-1}\bar{V}_{1}(\tau_{1}^{-1})^{\top}\big),$
where $\tau_{1}$ and $\bar{V}_{1}$ are defined as in Corollary 2.
Remark 3 (Why the parametric case does not require $o_{p}(N^{-1/2})$).
A correctly specified parametric model yields $\|\hat{b}-b^{*}\|_{2}=O_{p}(N^{-1/2})$, not $o_{p}(N^{-1/2})$. These are distinct. The parametric case does not need the $o_{p}(N^{-1/2})$ condition because the estimation error $r(O)=\dot{b}(O;\alpha_{0})^{\top}(\hat{\alpha}-\alpha_{0})$ is linear in a finite-dimensional $(\hat{\alpha}-\alpha_{0})$, which can be linearized via its estimating equation. The resulting $O_{p}(N^{-1/2})$ remainder is not negligible, but it is absorbed into the influence function (C.2) through the correction term $c_{\alpha}^{\top}A^{-1}\psi_{i}$. This parallels the treatment of propensity-score estimation via the $\kappa^{*}$ term in Corollary 1.
For a nonparametric $\hat{b}$, the estimation error $r(O)=\hat{b}(O)-b^{*}(O)$ is an infinite-dimensional object that cannot be linearized via a finite-dimensional parameter. The bias term $\mathbb{E}\{[\pi(O)\omega^{*}(O)-1]r(O)\}$ is a functional of the entire error function $r(\cdot)$, and there is no natural way to absorb it into a finite-dimensional influence function. Hence the rate condition (C.9) is genuinely needed.
Remark 4 (When does $c_{\alpha}=0$?).
The correction term $c_{\alpha}$ in (C.8) vanishes under the following conditions:
  1. (a)
    If $b(O;\alpha)=\alpha^{\top}\phi(O)$ is linear in $\alpha$ for some basis $\phi$, then $\dot{b}(O;\alpha_{0})=\phi(O)$. By the balancing constraint, $\mathbb{E}[(\delta\omega^{*}-1)\phi(O)]=0$, and by iterated expectations $\mathbb{E}[(\pi\omega^{*}-1)\phi]=\mathbb{E}[(\delta\omega^{*}-1)\phi]=0$, giving $c_{\alpha}=0$.
  2. (b)
    More generally, if $\dot{b}(O;\alpha_{0})\in\mathrm{span}\{b(O),g(\pi^{-1}(O;\varphi^{*}))\}$, then $c_{\alpha}=0$ by the calibration equations.
When $c_{\alpha}=0$, the parametric estimation of $b$ has no first-order effect on the asymptotic variance, and the oracle results in Corollary 2 hold exactly. This is analogous to the result $\kappa^{*}=0$ in Lemma 4.3, which states that PS estimation has no first-order effect when the OR model is correct.
Remark 5 (When is $\|\hat{b}-b^{*}\|_{2}=o_{p}(N^{-1/2})$ achievable?).
The $L_{2}$ convergence rate of nonparametric estimators depends on the smoothness of $b^{*}$ and the dimension of $O$. If $b^{*}(O)$ belongs to a Sobolev space $\mathcal{W}^{s,2}(\mathbb{R}^{d})$ with smoothness $s$ and $O\in\mathbb{R}^{d}$, the minimax $L_{2}$ rate is $N^{-s/(2s+d)}$, and condition (C.9) requires $s>d$, which is a strong restriction.
For additive models (GAM), if $b^{*}(O)=\sum_{j=1}^{d}b_{j}^{*}(O_{j})$ with each component in $\mathcal{W}^{s,2}(\mathbb{R})$, the effective dimension is $d_{\mathrm{eff}}=1$ and the rate becomes $N^{-s/(2s+1)}$. The condition then reduces to $s>1$, which is mild (twice differentiable suffices). This is relevant because the simulations in Section 7 use GAM as the prediction model.
C.4 Proof Sketch for Proposition C.1
The proof proceeds in three steps.
Step 1: Weights at $\lambda^{*}$ are independent of $\hat{b}$.  Under the correct PS model, Lemma 4.2 gives $\lambda^{*}=(0^{\top},1)^{\top}$, so the oracle calibration weights are
$\omega_{i}^{*}(\lambda^{*},s_{i}^{*})=g^{-1}\big(0^{\top}b_{i}^{*}+1\cdot g(\pi_{i}^{-1})\big)=g^{-1}(g(\pi_{i}^{-1}))=\pi_{i}^{-1}.$
Crucially, when we substitute $\hat{s}_{i}$ for $s_{i}^{*}$, the weights at $\lambda^{*}$ remain unchanged:
$\omega_{i}^{*}(\lambda^{*},\hat{s}_{i})=g^{-1}\big(0^{\top}\hat{b}_{i}+1\cdot g(\hat{\pi}_{i}^{-1})\big)=\hat{\pi}_{i}^{-1}.$ (C.10)
Since $\lambda_{1}^{*}=0$, the $\hat{b}_{i}$ component is annihilated, and the weights at the oracle $\lambda^{*}$ are exactly the IPW weights regardless of $\hat{b}$.
Step 2: Consistency of $\hat{\lambda}$ and its deviation from $\lambda^{*}$.  The estimated $\hat{\lambda}$ solves the gradient equation $\nabla\hat{\rho}_{G}(\lambda)=0$ with $\hat{s}_{i}$ replacing $s_{i}^{*}$:
$\frac{1}{N}\sum_{i=1}^{N}\big[\delta_{i}\,\omega_{i}^{*}(\lambda,\hat{s}_{i})\,\hat{s}_{i}-\hat{s}_{i}\big]=0.$ (C.11)
Evaluating the gradient at $\lambda^{*}$, the $\hat{b}-$component is
$\frac{1}{N}\sum_{k=1}^{K}\sum_{i\in\mathcal{I}^{(k)}}\Big(\frac{\delta_{i}}{\pi(O_{i};\varphi_{0})}-1\Big)\hat{b}_{i}^{(-k)}.$
Conditional on the training data $\mathcal{S}^{(-k)}$, the function $\hat{b}_{i}^{(-k)}$ is deterministic in $O_{i}$, and $\mathbb{E}[(\delta_{i}/\pi_{i}-1)\mid O_{i}]=0$ under the correct PS model. Therefore, within each fold:
$\mathbb{E}\bigg[\Big(\frac{\delta_{i}}{\pi_{i}}-1\Big)\hat{b}_{i}^{(-k)}\;\Big|\;\text{training}^{(-k)}\bigg]=0,$
so the gradient at $\lambda^{*}$ is a sum of conditionally centered terms with conditional variance
$\frac{1}{N}\mathbb{E}\bigg[\Big(\frac{1}{\pi_{i}}-1\Big)\big\|\hat{b}_{i}^{(-k)}\big\|^{2}\;\Big|\;\text{training}^{(-k)}\bigg]=O(1),$
which is bounded under (C.4) and positivity. By the CLT within each fold, the full gradient is $O_{p}(N^{-1/2})$, and the standard Taylor expansion gives
$\hat{\lambda}-\lambda^{*}=-\big[\mathbb{E}\{\nabla^{2}\rho_{G}(\lambda^{*})\}\big]^{-1}\nabla\hat{\rho}_{G}(\lambda^{*})+o_{p}(N^{-1/2}).$ (C.12)
Step 3: Linearization of $\hat{U}^{\omega}(\theta)$.  Following the proof of Theorem 4.1, expand around $\lambda^{*}$:
$\hat{U}^{\omega}(\theta)$ $=\frac{1}{N}\sum_{i=1}^{N}\delta_{i}\,\omega_{i}^{*}(\lambda^{*},\hat{s}_{i})\,U(\theta;z_{i})+\hat{\gamma}^{*}\cdot\nabla\hat{\rho}_{G}(\lambda^{*})+o_{p}(N^{-1/2})$
$=\frac{1}{N}\sum_{i=1}^{N}\frac{\delta_{i}}{\hat{\pi}_{i}}\,U(\theta;z_{i})+\hat{\gamma}^{*}\cdot\nabla\hat{\rho}_{G}(\lambda^{*})+o_{p}(N^{-1/2}),$ (C.13)
where the second equality uses (C.10). The first term is the standard IPW estimating equation, which is independent of $\hat{b}$. The second term is $O_{p}(N^{-1/2})$ by Step 2. The projection coefficient $\hat{\gamma}^{*}$ converges to $\gamma^{*}$ as defined in Theorem 4.1.
The key observation is that neither leading term depends on the quality of $\hat{b}$—only on its boundedness. The estimation error $r_{i}=\hat{b}_{i}-b_{i}^{*}$ enters through the gradient $\nabla\hat{\rho}_{G}(\lambda^{*})$, which has conditional mean zero under the correct PS model. Therefore, $r_{i}$ contributes only to the variance of $\nabla\hat{\rho}_{G}(\lambda^{*})$, not to its mean, and the overall contribution is absorbed into the $O_{p}(N^{-1/2})$ term.
Combining with the Taylor expansion of $\hat{U}^{\omega}$ around $\varphi^{*}$ (as in the proof of Corollary 1) yields the stated result with asymptotic variance $\tau_{1}^{-1}V_{1}(\tau_{1}^{-1})^{\top}$, identical to the oracle case. ∎
C.5 Proof Sketch for Proposition C.2 (Parametric OR)
When the OR model is correct and $\hat{b}$ is parametric, $\lambda_{1}^{*}\neq 0$ in general, and the estimation error in $\hat{b}$ directly affects the calibration weights. However, the parametric structure allows the estimation effect to be absorbed into the influence function.
Step 1: Taylor expansion of the remainder.  Write $r_{i}=b(O_{i};\hat{\alpha})-b(O_{i};\alpha_{0})$ and expand:
$r_{i}=\dot{b}(O_{i};\alpha_{0})^{\top}(\hat{\alpha}-\alpha_{0})+O_{p}(N^{-1}),$ (C.14)
where $\dot{b}(O;\alpha)=\partial b(O;\alpha)/\partial\alpha$. Decompose the estimating equation as in (C.18) below (Step 1 of Section C.6), obtaining the remainder
$R_{N}=\frac{1}{N}\sum_{i=1}^{N}(\delta_{i}\hat{\omega}_{i}-1)\,r_{i}.$ (C.15)
Decompose $\delta_{i}\hat{\omega}_{i}-1=\delta_{i}(\hat{\omega}_{i}-\omega_{i}^{*})+(\delta_{i}\omega_{i}^{*}-1)$, giving $R_{N}=R_{N,1}+R_{N,2}$ as in (C.20) below.
Step 2: $R_{N,1}$ is $o_{p}(N^{-1/2})$.  By Cauchy–Schwarz, $|R_{N,1}|\leq\|(\hat{\omega}-\omega^{*})\|_{2,n}\cdot\|r\|_{2,n}$. The first factor is $O_{p}(N^{-1/2})$ (Theorem 4.1) and the second is $O_{p}(N^{-1/2})$ from (C.14). Hence $R_{N,1}=O_{p}(N^{-1})=o_{p}(N^{-1/2})$.
Step 3: $R_{N,2}$ is $O_{p}(N^{-1/2})$ and absorbed into the influence function.  Substituting (C.14) into the critical term $R_{N,2}$:
$R_{N,2}$ $=\frac{1}{N}\sum_{i=1}^{N}(\delta_{i}\omega_{i}^{*}-1)\,\dot{b}(O_{i};\alpha_{0})^{\top}(\hat{\alpha}-\alpha_{0})+o_{p}(N^{-1/2})$
$=\left[\frac{1}{N}\sum_{i=1}^{N}(\delta_{i}\omega_{i}^{*}-1)\,\dot{b}(O_{i};\alpha_{0})\right]^{\top}(\hat{\alpha}-\alpha_{0})+o_{p}(N^{-1/2}).$ (C.16)
The bracketed term converges in probability to $c_{\alpha}$ as defined in (C.8). Substituting the linearization (C.5) of $\hat{\alpha}-\alpha_{0}$:
$R_{N,2}=-c_{\alpha}^{\top}A^{-1}\frac{1}{N}\sum_{i=1}^{N}\psi(z_{i},\delta_{i};\alpha_{0})+o_{p}(N^{-1/2}).$ (C.17)
This is not negligible—it is $O_{p}(N^{-1/2})$—but it is a sample average of i.i.d. terms that can be combined with the oracle influence function. Adding (C.17) to the oracle representation from Corollary 1 yields the modified influence function (C.2).
Step 4: Conditions for $c_{\alpha}=0$.  If $b(O;\alpha)$ is linear in $\alpha$, i.e., $b(O;\alpha)=\alpha^{\top}\phi(O)$, then $\dot{b}(O;\alpha_{0})=\phi(O)$. By the balancing constraint (3.3), $\mathbb{E}[\delta\omega^{*}\phi(O)]=\mathbb{E}[\phi(O)]$, so $\mathbb{E}[(\delta\omega^{*}-1)\phi(O)]=0$. Since $\mathbb{E}[(\pi\omega^{*}-1)\phi]=\mathbb{E}[\mathbb{E}[(\delta\omega^{*}-1)\phi\mid O]]=0$ by iterated expectations, we obtain $c_{\alpha}=0$, and the oracle variance $\bar{V}_{1}$ from Corollary 2 is recovered. ∎
C.6 Proof Sketch for Proposition C.3 (Nonparametric/ML OR)
When the OR model is correct but $\hat{b}$ is a nonparametric/ML estimator with cross-fitting, $\lambda_{1}^{*}\neq 0$ in general, and the estimation error in $\hat{b}$ directly affects the calibration weights. Unlike the parametric case, the infinite-dimensional estimation error cannot be absorbed into a finite-dimensional influence function.
Step 1: Decomposition of the estimating equation.  Using the balancing constraint $N^{-1}\sum\delta_{i}\hat{\omega}_{i}\hat{b}_{i}=N^{-1}\sum\hat{b}_{i}$, decompose:
$\hat{U}^{\omega}(\theta)$ $=\frac{1}{N}\sum_{i=1}^{N}\delta_{i}\hat{\omega}_{i}\,U(\theta;z_{i})$
$=\frac{1}{N}\sum_{i=1}^{N}b^{*}(O_{i})-R_{N}+\frac{1}{N}\sum_{i=1}^{N}\delta_{i}\hat{\omega}_{i}\big[U(\theta;z_{i})-b^{*}(O_{i})\big],$ (C.18)
where the remainder from replacing $b^{*}$ by $\hat{b}$ in the balancing constraint is
$R_{N}=\frac{1}{N}\sum_{i=1}^{N}(\delta_{i}\hat{\omega}_{i}-1)\,r_{i},\qquad r_{i}=\hat{b}_{i}-b^{*}(O_{i}).$ (C.19)
Step 2: Analysis of the remainder $R_{N}$.  Decompose $\delta_{i}\hat{\omega}_{i}-1=\delta_{i}(\hat{\omega}_{i}-\omega_{i}^{*})+(\delta_{i}\omega_{i}^{*}-1)$, giving
$R_{N}=\underbrace{\frac{1}{N}\sum_{i=1}^{N}\delta_{i}(\hat{\omega}_{i}-\omega_{i}^{*})\,r_{i}}_{R_{N,1}}+\underbrace{\frac{1}{N}\sum_{i=1}^{N}(\delta_{i}\omega_{i}^{*}-1)\,r_{i}}_{R_{N,2}}.$ (C.20)
Term $R_{N,1}$ (product of estimation errors).  By the Cauchy–Schwarz inequality and cross-fitting independence:
$|R_{N,1}|\leq\bigg(\frac{1}{N}\sum_{i=1}^{N}\delta_{i}(\hat{\omega}_{i}-\omega_{i}^{*})^{2}\bigg)^{1/2}\bigg(\frac{1}{N}\sum_{i=1}^{N}\delta_{i}\,r_{i}^{2}\bigg)^{1/2}.$
The first factor is $O_{p}(N^{-1/2})$ from the linearization of $\hat{\omega}$ (Theorem 4.1). Under (C.4), the second factor is $o_{p}(1)$. Hence $R_{N,1}=o_{p}(N^{-1/2})$ under $L_{2}$-consistency alone.
Term $R_{N,2}$ (direct remainder).  This is the critical term. Within fold $k$, conditional on training data $\mathcal{S}^{(-k)}$:
$\mathbb{E}\big[(\delta_{i}\omega_{i}^{*}-1)\,r_{i}(O_{i})\;\big|\;\text{training}^{(-k)}\big]=\mathbb{E}\big\{[\pi(O)\,\omega^{*}(O;\lambda^{*},\varphi^{*})-1]\,r(O)\big\},$ (C.21)
which is generally nonzero when $r(O)\neq 0$. By the calibration equation, $\mathbb{E}[(\delta\omega^{*}-1)s^{*}]=0$, so $\mathbb{E}[(\delta\omega^{*}-1)b^{*}(O)]=0$ since $b^{*}(O)$ is a subvector of $s^{*}$. However, $r(O)=\hat{b}(O)-b^{*}(O)\notin\mathrm{span}\{s^{*}\}$ in general, so the conditional mean (C.21) does not vanish.
By the Cauchy–Schwarz inequality:
$\big|\mathbb{E}\{[\pi\omega^{*}-1]\,r(O)\}\big|\leq\big(\mathbb{E}[(\pi\omega^{*}-1)^{2}]\big)^{1/2}\,\|r\|_{2}=C\cdot\|r\|_{2}.$
The conditional variance within each fold is $O(\|r\|_{2}^{2}/N)$. Combining across folds:
$R_{N,2}=O(\|r\|_{2})+O_{p}(N^{-1/2}\|r\|_{2}).$ (C.22)
For $R_{N,2}=o_{p}(N^{-1/2})$, the dominant condition is $\|r\|_{2}=o_{p}(N^{-1/2})$, which is precisely condition (C.9).
Why the parametric trick of Proposition C.2 fails here.  For a nonparametric $\hat{b}$, the estimation error $r(O)=\hat{b}(O)-b^{*}(O)$ is an infinite-dimensional object with no finite-dimensional parametric representation. The bias term $\mathbb{E}\{[\pi\omega^{*}-1]r(O)\}$ is a functional of the entire error function $r(\cdot)$, and cannot be factored as $c^{\top}(\hat{\alpha}-\alpha_{0})$ for any finite-dimensional $c$ and $\hat{\alpha}$. Consequently, there is no estimating equation through which to absorb the $O(\|r\|_{2})$ bias into the influence function, and the rate condition (C.9) is necessary. ∎
C.7 Practical Implications
These results suggest the following practical guidance:
  1. 1.
    When the practitioner has confidence in the PS model (e.g., in randomized experiments or well-understood selection mechanisms), cross-fitted ML predictions of any quality can be used for $\hat{b}$ to gain efficiency without risking consistency.
  2. 2.
    When the PS model is suspect and the practitioner relies on the OR pathway with a parametric outcome model, no rate condition is needed. The estimation effect of $\hat{\alpha}$ is automatically accounted for in the influence function, and contributes zero additional asymptotic variance when $\dot{b}(O;\alpha_{0})$ lies in the calibration space.
  3. 3.
    When the PS model is suspect and a nonparametric/ML predictor is used for $\hat{b}$, the ML predictor $\hat{m}^{(-k)}$ should converge at a rate faster than $N^{-1/2}$ in $L_{2}$-norm. For fully nonparametric methods in $d$ dimensions, this requires smoothness $s>d$; for additive models (GAM), the much milder condition $s>1$ suffices.
  4. 4.
    The use of GAM in the simulations (Section 7) is well-motivated from this perspective: the additive structure ensures condition (C.9) holds under mild smoothness, avoiding the curse of dimensionality inherent in general nonparametric estimation.
C.8 Summary
The following table summarizes the rate conditions for the cross-fitted GEC estimator:
Scenario Condition on $\hat{b}$ Effect on asymptotics
PS correct, $\hat{b}$ arbitrary $\|r\|_{2}=o_{p}(1)$ Oracle variance $V_{1}$
OR correct, $\hat{b}$ parametric $O_{p}(N^{-1/2})$ (automatic) Modified IF via $c_{\alpha}$;
$c_{\alpha}=0$ if $\dot{b}\in\mathrm{span}\{s^{*}\}$
OR correct, $\hat{b}$ nonparametric $\|r\|_{2}=o_{p}(N^{-1/2})$ Oracle variance $\bar{V}_{1}$
Both correct $\|r\|_{2}=o_{p}(1)$ Semiparametric efficient